### v1

Abstract    Smoluchowski-Kramers approximation in rough path topology with state-dependent damping is explored. The second-order Langevin equation has a form of fast-slow system after suitable change-of-variable, and then its solution is lifted as a rough path in a natural manner. Moment estimates of both the original path and the lift are given, followed by which, averaging technique and convergence theorem in rough path topology are used to pass the limit.
Keywords    Smoluchowski–Kramers approximation, Averaging, Singular pertubation, Rough path theory
2020 Mathematics Subject Classification   60L90, 60H10
1 Introduction
Consider the state-dependent second-order Langevin equation
$\epsilon^{2}\ddot{X}^{\epsilon}_{t}+M(X^{\epsilon}_{t})\dot{X}^{\epsilon}_{t}=F(X^{\epsilon}_{t})+\dot{W}_{t},\kern 5.0ptX^{\epsilon}_{0}=0,\kern 5.0pt\dot{X}^{\epsilon}_{0}=0,$ (1.1)
defined on a complete stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{0\leq t\leq T},P),$ where $0<T<\infty$ is fixed, $M$ is a matrix-valued function and $F$ is a non-linearity, detailed in Section 2. We are interested in the limit of $X^{\epsilon}$ as $\epsilon\to 0,$ which is called Smoluchowski–Kramers (SK) approximation. There is plenty of research for the case $M$ is a constant [2, 3, 33, 16, 30, 7], and in this case, it is not surprising that $X^{\epsilon}\to X$ in some sense with
$M\dot{X}_{t}=F(X_{t})+\dot{W}_{t},\kern 5.0ptX_{0}=0,$
which can be obtained formally by taking $\epsilon=0$ in (1.1). However, when $M$ is not a constant, that is, it is state-dependent, the limit cannot be obtained by simply dropping the small mass term. In fact, under certain assumptions, Hottovy et al.[12] showed that
$\lim_{\epsilon\rightarrow 0}\mathbb{E}\sup\limits_{0\leq t\leq T}|X^{\epsilon}_{t}-X_{t}|^{2}=0,$
where
$dX_{t}=S(X_{t})dt+M^{-1}(X_{t})F(X_{t})dt+M^{-1}(X_{t})dW_{t},\kern 5.0ptX_{0}=0,$ (1.2)
and
$S_{j}(x)=\sum_{k,l}(\partial_{x_{l}}M^{-1})_{jk}(x)J_{kl}(x),$
with $J$ the solution of Lyapunov equation
$M(x)J(x)+J(x)M^{\top}(x)=\mathrm{id}.$ (1.3)
Since then, there are fruitful work on SK approximation with state-dependent damping. For example, Wang et al.[27] explore the limit by averaging on the Fokker-Planck equation associated with the Langevin equation. Cerrai and Xi[4] consider the infinite dimensional case. For more research in SK approximation in state dependent case, we refer to [5, 11, 17, 28].
SK approximation is naturally linked to stochastic averaging, initiated by Khaminskii [13]. Numerous work concerning about stochastic averaging appeared in recent years, see e.g. [14, 29, 32, 31]. There is also research about averaging principle in rough path framework, see for example in [15, 25]. In fact, by suitable change-of-variable, (1.1) is transformed to (2.3), which has a form of fast-slow system [6], making it possible to find its limit by averaging method. A Poisson equation technique[21, 22, 23] is employed to show the convergence.
Rough path theory is initially proposed by Lyons[19]. Although a deterministic theory, it is proved a powerful tool in stochastic analysis. See monographs [18, 20, 9, 8] for systematic display of rough path theory. In this paper, we aim to find the small mass limit in rough path topology. Friz et al [7] identify the limit in rough path topology in the case that $M$ is a constant matrix and $F\equiv 0$. They found that, the limit of the "level-2" rough path, converges to a non-trivial limit. In particular, when the dimension is $1$, the Itô lift of the original path, converges to the Stratonovich lift of the limit path, which is reminiscent of Wong–Zakai approximation [26]. They also gave a conjecture for the case when $M$ is not constant, without detailed proof. In this paper, the system with both state-dependent friction and non-linearity is explored. We identify the limit in the rough path topology, and give rigorous proof. This also generalizes the result in [12], in the sense that in our main theorem, the "level-1" path converges in $p$-th moment of $\alpha$-Hölder norm for all $1\leq p<\infty$.
The rest of this paper is organized as follows. In Section 2, we recall some basic rough path theory, list some notations, and present our main result. In Section 3, we give moment estimate in rough path norm. In Section 4, we display an averaging principle for specific fast-slow system, which is used to pass the limit. After these preparation, we prove the main theorem in Section 5.
2 Preliminary and Main Result
We recall some basic definitions of rough path theory [8, e.g.]. Let $X:[0,T]\to\mathbb{R}^{d},$ $\mathbb{X}:\Delta_{[0,T]}\to\mathbb{R}^{d\times d}$, where
$\Delta_{[0,T]}:=\{(s,t)\in[0,T]^{2}:s<t\}.$
Given $\frac{1}{3}<\alpha\leq\frac{1}{2},$ a map $\mathbf{X}=(X,\mathbb{X})$ is called an $\alpha$-Hölder rough path over $\mathbb{R}^{d}$, if the following conditions hold:
(i) $X||_{\alpha}:=\sup_{0\leq s\leq t}\limits\frac{|X_{t}-X_{s}|}{|t-s|^{\alpha}}<\infty$
(ii) $\mathbb{X}||_{2\alpha}:=\sup_{0\leq s\leq t}\limits\frac{|\mathbb{X}_{s,t}|}{|t-s|^{2\alpha}}<\infty$
(iii) (Chen relation) $\mathbb{X}_{s,u}=\mathbb{X}_{s,t}+\mathbb{X}_{t,u}+X_{s,t}\otimes X_{t,u},\kern 5.0pt\forall 0\leq s\leq t\leq u\leq T.$
The set of all $\alpha$-Hölder rough path over $\mathbb{R}^{d}$ is denoted as $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d}).$ Given $\mathbf{X},\mathbf{Y}\in\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d}),$ the $\alpha$-Hölder rough path metric is defined as
[…]
Let $V^{\epsilon}_{t}=\dot{X}^{\epsilon}_{t}$ and $P^{\epsilon}_{t}=\epsilon^{2}V^{\epsilon}_{t}$ be velocity and momentum, respectively, then (1.1) is equivalent to
$\begin{cases}dX^{\epsilon}_{t}=V^{\epsilon}_{t}dt,\kern 5.0ptX^{\epsilon}_{0}=0,\\ dV^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})V^{\epsilon}_{t}dt+\frac{1}{\epsilon^{2}}F(X^{\epsilon}_{t})dt+\frac{1}{\epsilon^{2}}dW_{t},\kern 5.0ptV^{\epsilon}_{0}=0,\end{cases}$ (2.1)
and
$\begin{cases}dX^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}P^{\epsilon}_{t}dt,\kern 5.0ptX^{\epsilon}_{0}=0,\\ dP^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})P^{\epsilon}_{t}dt+F(X^{\epsilon}_{t})dt+dW_{t},\kern 5.0ptP^{\epsilon}_{0}=0.\end{cases}$ (2.2)
Here we also introduce $Y^{\epsilon}_{t}:=\frac{1}{\epsilon}P^{\epsilon}_{t}$ [7, Page 7944], so that
$\begin{cases}dX^{\epsilon}_{t}=\frac{1}{\epsilon}Y^{\epsilon}_{t}dt,\kern 5.0ptX^{\epsilon}_{0}=0,\\ dY^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})Y^{\epsilon}_{t}dt+\frac{1}{\epsilon}F(X^{\epsilon}_{t})dt+\frac{1}{\epsilon}dW_{t},\kern 5.0ptY^{\epsilon}_{0}=0.\end{cases}$ (2.3)
For each $0<\epsilon\leq 1,$ define the lift of $X^{\epsilon}$ by
$\mathbb{X}^{\epsilon}_{s,t}:=\int_{s}^{t}X^{\epsilon}_{s,u}\otimes dX^{\epsilon}_{u},\quad(s,t)\in\Delta_{[0,T]},$
where the stochastic integral is Itô integral, and set $\mathbf{X}^{\epsilon}=(X^{\epsilon},\mathbb{X}^{\epsilon}).$ It is straightforward to check that $\mathbf{X}^{\epsilon}$ satisfies Chen relation. Moreover, by Kolmogorov criterion for rough paths [8, Theorem 3.1] and Proposition 3.3 below, $\mathbf{X}^{\epsilon}$ lies in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$ for all $\frac{1}{3}<\alpha<\frac{1}{2}.$
We impose the following assumptions:
($\mathbf{A1}$) There exists a constant $\lambda>0$, such that the smallest eigenvalue $\lambda_{1}$ of $\frac{M+M^{\top}}{2}$ satisfies that
$\lambda_{1}(x)\geq\lambda,\quad x\in\mathbb{R}^{d}.$
($\mathbf{A2}$) $M\in C^{2}(\mathbb{R}^{d};\mathbb{R}^{d\times d})$, and $M^{-1}\in C^{3}(\mathbb{R}^{d};\mathbb{R}^{d\times d})$.
($\mathbf{A3}$) $M$, $M^{-1}$ and $\partial_{x_{j}}M^{-1}_{ik}$ are globally Lipschitz, and $M^{-1}$ is bounded, $1\leq i,j,k\leq d$.
($\mathbf{A4}$) $F$ is globally Lipschitz and bounded.
Remark 2.1.
From ($\mathbf{A1}$), one finds that
$y^{\top}M(x)y\geq\lambda|y|^{2},\quad(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d},$
which is essential in moment estimates. ($\mathbf{A2}$) is used to obtain regularity of the solution of Poisson equation in Section 4. Well-posedness of (1.2) and (2.3) are guaranteed by ($\mathbf{A3}$) and ($\mathbf{A4}$).
Now we state the main result.
Theorem 2.2.
Let $X$ be the solution of (1.2). Define
$\mathbb{X}_{s,t}:=\int_{s}^{t}X_{s,u}\otimes\circ dX_{u}+\frac{1}{2}\int_{s}^{t}J(X_{u})(M^{-1})^{\top}(X_{u})-M^{-1}(X_{u})J(X_{u})du,$ (2.4)
where $J$ is the solution of (1.3), and $\otimes\circ$ denotes the Stratonovich integral. Under assumptions $(\mathbf{A1})$$(\mathbf{A4})$, for each $\frac{1}{3}<\alpha<\frac{1}{2}$ and $1\leq p<\infty,$
$\lim\limits_{\epsilon\to 0}\mathbb{E}[\rho_{\alpha}(\mathbf{X}^{\epsilon},\mathbf{X})^{p}]=0,$
where $\mathbf{X}=(X,\mathbb{X}).$
Remark 2.3.
By Kolmogorov criterion for rough paths again, one verifies that $\mathbf{X}$ lies in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$ for all $\frac{1}{3}<\alpha<\frac{1}{2}.$
We end this section with a collection of frequently used notation.
Notation
$|\cdot|$: Euclidean norm of either a vector or a matrix.
$A^{\top}$: Transpose of a matrix or a vector $A$.
$\mathrm{id}$: Identity matrix.
$(x,y)$: Inner product of two vectors $x$ and $y$.
$Df,D^{2}f$: Gradient and Hessian of a function $f:\mathbb{R}^{d}\to\mathbb{R}$.
$C^{k}(U;V)$: The set consists of $k$-th continuously differentiable functions from $U$ to $V$, where $k\in\mathbb{N},$ and $U$, $V$ are Banach spaces.
$a\lesssim b$: There exists a constant $C>0$ such that $a\leq Cb.$ Throughout the paper, $C$ never depends on $\epsilon$. Additionally, in the proof of Proposition 3.3, it neither depends on $s$ nor $t$.
3 Moment Estimates
In this section, we give several moment estimate for system (2.3). We begin with bounding the supremum norm.
Proposition 3.1.
Under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$), For each $1\leq p<\infty,$
$\sup\limits_{0<\epsilon\leq 1}\sup\limits_{0\leq t\leq T}\mathbb{E}|Y^{\epsilon}_{t}|^{p}<\infty,\kern 5.0pt\sup\limits_{0<\epsilon\leq 1}\mathbb{E}\sup\limits_{0\leq t\leq T}|X^{\epsilon}_{t}|^{p}<\infty.$
Proof.
We firstly show the boundedness for $Y^{\epsilon}$. For each $p\geq 2,$ set $\phi(y)=|y|^{p},$ then
$D\phi(y)=p|y|^{p-2}y,\quad D^{2}\phi(y)=p(p-2)|y|^{p-4}yy^{\top}+p|y|^{p-2}\mathrm{id}.$
Apply Itô formula to $\phi(Y^{\epsilon}_{t})$ and use ($\mathbf{A1}$) and ($\mathbf{A4}$) to find that
$|Y^{\epsilon}_{t}|^{p}$ (3.1)
$=$ $\int_{0}^{t}[D\phi(Y^{\epsilon}_{s})]^{\top}dY^{\epsilon}_{s}+\frac{1}{2}\int_{0}^{t}\mathrm{Tr}\Big(\frac{1}{\epsilon^{2}}D^{2}\phi(Y^{\epsilon}_{s})\Big)ds$
$=$ $-\frac{p}{\epsilon^{2}}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-2}(Y^{\epsilon}_{s})^{\top}M(X^{\epsilon}_{s})Y^{\epsilon}_{s}ds+\frac{p}{\epsilon}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-2}(Y^{\epsilon}_{s})^{\top}F(X^{\epsilon}_{s})ds+H^{\epsilon}_{t}$
$+\frac{1}{2\epsilon^{2}}\int_{0}^{t}\mathrm{Tr}[p(p-2)|Y^{\epsilon}_{s}|^{p-4}Y^{\epsilon}_{s}(Y^{\epsilon}_{s})^{\top}+p|Y^{\epsilon}_{s}|^{p-2}\mathrm{id}]$
$\lesssim$ $-\frac{1}{\epsilon^{2}}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p}ds+\frac{1}{\epsilon}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-1}ds+H^{\epsilon}_{t}+\frac{1}{\epsilon^{2}}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-2}ds,$
where $H^{\epsilon}$ is a martingale. Taking expectation and applying Hölder inequality,
$\mathbb{E}|Y^{\epsilon}_{t}|^{p}\lesssim-\frac{1}{\epsilon^{2}}\int_{0}^{t}\mathbb{E}|Y^{\epsilon}_{s}|^{p}ds+\frac{1}{\epsilon^{2}}\int_{0}^{t}\mathbb{E}|Y^{\epsilon}_{s}|^{p-2}ds.$ (3.2)
First for $p=2$, by the comparison principle,
$\sup\limits_{0<\epsilon\leq 1}\sup\limits_{0\leq t\leq T}\mathbb{E}|Y^{\epsilon}_{t}|^{2}<\infty.$
By (3.2) and induction,
$\sup\limits_{0<\epsilon\leq 1}\sup\limits_{0\leq t\leq T}\mathbb{E}|Y^{\epsilon}_{t}|^{p}<\infty.$
for all even $p$, and hence for all $1\leq p<\infty.$
Next, we show the boundedness for $X^{\epsilon}$. Consider the following random ODE in $\mathbb{R}^{d\times d}$
$\dot{\phi}^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})\phi^{\epsilon}_{t},\kern 5.0pt\phi^{\epsilon}_{0}=\mathrm{id}.$
For each $r\in[0,T],$ set $z^{\epsilon,r}_{s}=\phi^{\epsilon}_{s}(\phi^{\epsilon}_{r})^{-1},\kern 5.0pts\in[r,T].$ It is showed in [27, Page 6] that, under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$),
$X^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{0}^{s}z^{\epsilon,r}_{s}dW_{r}ds+\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{0}^{s}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})drds.$ (3.3)
and that
$|z^{\epsilon,r}_{s}|\leq e^{-\epsilon^{-2}\lambda(s-r)},\quad 0\leq r\leq s\leq T.$ (3.4)
For notation simplicity, we always assume that $\lambda=1$. Then, from (3.4)
$\Big|\int_{r}^{t}z^{\epsilon,r}_{s}ds\Big|\leq\int_{r}^{t}|z^{\epsilon,r}_{s}|ds\leq\epsilon^{2}.$ (3.5)
Notice that by stochastic Fubini theorem, we have
$X^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}dsdW_{r}+\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})dsdr.$ (3.6)
An application of Burkholder–Davis–Gundy inequality and (3.5) yield
$\frac{1}{\epsilon^{2p}}\mathbb{E}\sup\limits_{0\leq t\leq T}\Big|\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}dsdW_{r}\Big|^{p}$ (3.7)
$\lesssim$ $\frac{1}{\epsilon^{2p}}\mathbb{E}\Big(\int_{0}^{T}\Big|\int_{r}^{t}z^{\epsilon,r}_{s}ds\Big|^{2}dr\Big)^{p/2}$
$\lesssim$ $1.$
By ($\mathbf{A4}$) and (3.5), for each $t\in[0,T],$
$\frac{1}{\epsilon^{2p}}\Big|\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})dsdr\Big|^{p}$ (3.8)
$\leq$ $\frac{1}{\epsilon^{2p}}\Big(\int_{0}^{t}\int_{r}^{t}|z^{\epsilon,r}_{s}|dsdr\Big)^{p}$
$\leq$ $T^{p}.$
Taking supremum and expectation,
$\frac{1}{\epsilon^{2p}}\mathbb{E}\sup\limits_{0\leq t\leq T}\Big|\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})dsdr\Big|^{p}\lesssim 1.$ (3.9)
The result follows from combining (3.6), (3.7) and (3.9). ∎
We need the following elementary inequality, which is a special case of Minkowski integral inequality[10, Page 13]. For completeness, we give its proof.
Lemma 3.2.
Let $G$ be a random process on $[0,T],$ $a,b]\subset[0,T],$ $q\in\mathbb{N},$ then
$\Big[\mathbb{E}\Big(\int_{a}^{b}|G(r)|^{2}dr\Big)^{q}\Big]^{1/q}\leq\int_{a}^{b}\Big(\mathbb{E}|G(r)|^{2q}\Big)^{1/q}dr.$
Proof.
Applying Hölder inequality,
$\mathbb{E}\Big(\int_{a}^{b}|G(r)|^{2}dr\Big)^{q}$
$=$ $\mathbb{E}\int_{[a,b]^{q}}|G(r_{1})|^{2}\cdots|G(r_{q})|^{2}dr_{1}\cdots dr_{q}$
$=$ $\int_{[a,b]^{q}}\mathbb{E}(|G(r_{1})|^{2}\cdots|G(r_{q})|^{2})dr_{1}\cdots dr_{q}$
$\leq$ $\int_{[a,b]^{q}}(\mathbb{E}|G(r_{1})|^{2q})^{1/q}\cdots(\mathbb{E}|G(r_{q})|^{2q})^{1/q}dr_{1}\cdots dr_{q}$
$=$ $\Big(\int_{a}^{b}(\mathbb{E}|G(r)|^{2q})^{1/q}dr\Big)^{q}.$
The next proposition concerns about Hölder norm of the path $X^{\epsilon}$ and its lift.
Proposition 3.3.
Under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$), there is a constant $C$ independent of $\epsilon$, such that
(i) For each $1\leq p<\infty,$
$\mathbb{E}|X^{\epsilon}_{t}-X^{\epsilon}_{s}|^{p}\leq C|t-s|^{p/2},\quad 0\leq s\leq t\leq T.$
(ii) For each $1\leq p<\infty,$
$\mathbb{E}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}\otimes dX^{\epsilon}_{u}\Big|^{p}\leq C|t-s|^{p},\quad 0\leq s\leq t\leq T.$
Proof.
In view of Jensen inequality, it is sufficient to prove the proposition for all even $p.$ We first prove (i). From (3.3) and stochastic Fubini theorem,
$X^{\epsilon}_{t}-X^{\epsilon}_{s}$ (3.10)
$=$ $\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}z^{\epsilon,r}_{u}dW_{r}du+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})drdu$
$=$ $\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}dudW_{r}+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big]$
$+\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big].$
Now we deal with four terms above respectively. By Burkholder–Davis–Gundy inequality,
$\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}\lesssim\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}.$ (3.11)
From (3.4) and Hölder inequality
$\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}$
$\leq$ $\Big|\int_{s}^{t}e^{-\epsilon^{-2}(u-r)}du\Big|^{2}$
$\leq$ $(t-s)\int_{s}^{t}e^{-2\epsilon^{-2}(u-r)}du.$
Therefore,
$\int_{0}^{s}\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}dr\lesssim(t-s)\int_{0}^{s}\int_{s}^{t}e^{-2\epsilon^{-2}(u-r)}dudr\lesssim\epsilon^{4}(t-s).$ (3.12)
Substituting the above inequality into (3.11),
$\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}\lesssim\epsilon^{2p}|t-s|^{p/2}.$ (3.13)
By Burkholder–Davis–Gundy inequality and (3.4),
$\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}$ (3.14)
$\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}z^{\epsilon,r}_{u}du\Big|^{2}\Big)^{p/2}$
$\leq$ $\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}e^{-\epsilon^{-2}(u-r)}du\Big|^{2}dr\Big)^{p/2}$
$\leq$ $\Big(\int_{s}^{t}\epsilon^{4}dr\Big)^{p/2}$
$\leq$ $\epsilon^{2p}(t-s)^{p/2}.$
On the other hand, by ($\mathbf{A4}$), Hölder inequality and (3.12),
$\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.15)
$\lesssim$ $\mathbb{E}\Big(\int_{0}^{s}\int_{s}^{t}|z^{\epsilon,r}_{u}|dudr\Big)^{p}$
$=$ $\mathbb{E}\Big\{\Big[\Big(\int_{0}^{s}\int_{s}^{t}|z^{\epsilon,r}_{u}|dudr\Big)^{2}\Big]^{p/2}\Big\}$
$\lesssim$ $\mathbb{E}\Big\{\Big[\int_{0}^{s}\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big]^{p/2}\Big\}$
$\lesssim$ $\epsilon^{2p}|t-s|^{p/2}.$
Lastly,
$\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.16)
$\lesssim$ $\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}e^{-\epsilon^{-2}(u-r)}dudr\Big|^{p}$
$\leq$ $\Big|\int_{s}^{t}\epsilon^{2}dr\Big|^{p}$
$\leq$ $\epsilon^{2p}(t-s)^{p}.$
Substituting (3.13)–(3.16) into (3.10),
$\mathbb{E}|X^{\epsilon}_{t}-X^{\epsilon}_{s}|^{p}\lesssim|t-s|^{p/2}.$ (3.17)
We turn to prove (ii). By triangular inequality, it is sufficient to consider the $(i,j)$-entry $\int_{s}^{t}X^{\epsilon,i}_{s,u}dX^{\epsilon,j}_{u}$ of $\int_{s}^{t}X^{\epsilon}_{s,u}\otimes dX^{\epsilon}_{u}.$ But from (3.3),
$dX^{\epsilon,j}_{t}=\frac{1}{\epsilon^{2}}\sum_{k}\int_{0}^{t}z^{\epsilon,r}_{t}(j,k)dB^{k}_{r}dt+\frac{1}{\epsilon^{2}}\sum_{k}\int_{0}^{t}z^{\epsilon,r}_{t}(j,k)F_{k}(X^{\epsilon}_{r})drdt,$
where $z^{\epsilon,r}_{t}(j,k)$ is the ($j,k$)-entry of $z^{\epsilon,r}_{t}$, so
$\int_{s}^{t}X^{\epsilon,i}_{s,u}dX^{\epsilon,j}_{u}=\frac{1}{\epsilon^{2}}\sum_{k}\int_{s}^{t}\int_{0}^{u}X^{\epsilon,i}_{s,u}z^{\epsilon,r}_{u}(j,k)dB^{k}_{r}du+\frac{1}{\epsilon^{2}}\sum_{k}\int_{s}^{t}\int_{0}^{u}X^{\epsilon,i}_{s,u}z^{\epsilon,r}_{u}(j,k)F_{k}(X^{\epsilon}_{r})drdu.$
Without loss of generality, we assume that the system is one-dimensional, that is, we write
$\int_{s}^{t}X^{\epsilon}_{s,u}dX^{\epsilon}_{u}=\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dW_{r}du+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})drdu.$ (3.18)
By stochastic Fubini theorem,
$\int_{s}^{t}X^{\epsilon}_{s,u}dX^{\epsilon}_{u}$ (3.19)
$=$ $\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}\Big]$
$+\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big]$
$=:$ $I+J.$
We firstly estimate $I$. Using Burkholder–Davis–Gundy inequality,
$\mathbb{E}|I|^{p}$ (3.20)
$\lesssim$ $\frac{1}{\epsilon^{2p}}\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}+\frac{1}{\epsilon^{2p}}\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}$
$\lesssim$ $\frac{1}{\epsilon^{2p}}\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}+\frac{1}{\epsilon^{2p}}\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}.$
Applying Lemma 3.2 to $G(r)=\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du$ and $q=p/2,$
$\Big\{\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}\leq\int_{0}^{s}\Big(\mathbb{E}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{2/p}dr.$ (3.21)
Using Hölder inequality and taking $p$-th power,
$\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\leq\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}|^{2}du\Big)^{p/2}\Big(\int_{s}^{t}|z^{\epsilon,r}_{u}|^{2}\Big)^{p/2}.$ (3.22)
Applying Lemma 3.2 to $G(u)=|X^{\epsilon}_{s,u}|$ and $q=p/2,$ and then using (3.17),
$\Big\{\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}|^{2}du\Big)^{p/2}\Big\}^{2/p}$
$\leq$ $\int_{s}^{t}(\mathbb{E}|X^{\epsilon}_{s,u}|^{p})^{2/p}du$
$\lesssim$ $\int_{s}^{t}|u-s|du$
$=$ $\frac{(t-s)^{2}}{2}.$
Taking ($p/2$)-th power,
$\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}|^{2}du\Big)^{p/2}\lesssim|t-s|^{p}.$ (3.23)
On the other hand, from (3.4) we have
$\int_{s}^{t}|z^{\epsilon,r}_{u}|^{2}du\leq\int_{s}^{t}e^{-2\epsilon^{-2}(u-r)}du\leq\frac{\epsilon^{2}}{2}e^{-2\epsilon^{-2}(s-r)}.$
Taking $(p/2)$-th power,
$\Big(\int_{s}^{t}|z^{\epsilon,r}_{u}|^{2}du\Big)^{p/2}\lesssim\epsilon^{p}e^{-p\epsilon^{-2}(s-r)}.$ (3.24)
Combining (3.22), (3.23) and (3.24),
$\mathbb{E}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\leq\epsilon^{p}e^{-p\epsilon^{-2}(s-r)}|t-s|^{p}.$
Taking ($2/p$)-th power and substituting into (3.21),
$\Big\{\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}$
$\lesssim$ $\int_{0}^{s}\epsilon^{2}e^{-2\epsilon^{-2}(s-r)}|t-s|^{2}dr$
$=$ $\epsilon^{2}(t-s)^{2}\int_{0}^{s}e^{-2\epsilon^{-2}(s-r)}dr$
$\leq$ $\epsilon^{4}(t-s)^{2}.$
Taking ($p/2$)-th power on the above inequality,
$\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\lesssim\epsilon^{2p}(t-s)^{p}.$ (3.25)
We deal with the other term in a similar fashion. Applying Lemma 3.2 to $G(r)=\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du$ and $q=p/2,$
$\Big\{\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}\leq\int_{s}^{t}\Big(\mathbb{E}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{2/p}dr.$ (3.26)
For fixed $r\in[s,t],$ using Lemma 3.2 to $G(u)=\sqrt{|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|},u\in[r,t]$ and $q=p$,
$\Big(\mathbb{E}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{1/p}$ (3.27)
$\leq$ $\Big(\mathbb{E}\Big[\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|du\Big]^{p}\Big)^{1/p}$
$\leq$ $\int_{r}^{t}\Big(\mathbb{E}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|^{p})^{1/p}du.$
By (3.4) and (3.17),
$\mathbb{E}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|^{p}\leq e^{-p\epsilon^{-2}(u-r)}\mathbb{E}|X^{\epsilon}_{s,u}|^{p}\leq e^{-p\epsilon^{-2}(u-r)}|u-s|^{p/2}.$
Taking $(1/p)$-th power and substituting it into (3.27),
$\Big(\mathbb{E}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{1/p}\leq\int_{r}^{t}e^{-\epsilon^{-2}(u-r)}|u-s|^{1/2}du\leq\epsilon^{2}|t-s|^{1/2}.$ (3.28)
Taking square and substituting it into (3.26),
$\Big\{\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}\leq\int_{s}^{t}\epsilon^{4}|t-s|dr=\epsilon^{4}|t-s|^{2}.$ (3.29)
Taking $(p/2)$-th power,
$\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\leq\epsilon^{2p}|t-s|^{p}.$ (3.30)
Combining (3.20), (3.25) and (3.30),
$\mathbb{E}|I|^{p}\lesssim|t-s|^{p}.$ (3.31)
It remains to deal with $J$. Writing
$\Big|\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.32)
$\lesssim$ $\Big(\int_{0}^{s}\int_{s}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{p}$
$=$ $\Big[\Big(\int_{0}^{s}\int_{s}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{2}\Big]^{p/2}$
$\lesssim$ $\Big[\int_{0}^{s}\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|du\Big)^{2}dr\Big]^{p/2},$
and
$\Big|\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.33)
$\lesssim$ $\Big(\int_{s}^{t}\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{p}$
$=$ $\Big[\Big(\int_{s}^{t}\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{2}\Big]^{p/2}$
$\lesssim$ $\Big[\int_{s}^{t}\Big(\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|du\Big)^{2}dr\Big]^{p/2},$
so that we notice that the estimate of $J$ is identical to $I$ in view of (3.20), from which the proof is ended. ∎
An immediate consequence of Proposition 3.3 and Kolmogorov criterion for rough paths [8, Theorem 3.1] is the following corollary.
Corollary 3.4.
Under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$), for all $1\leq p<\infty$ and $\frac{1}{3}<\alpha<\frac{1}{2},$
[…]
Remark 3.5.
From Corollary 3.4, we see that $(X^{\epsilon})_{0<\epsilon\leq 1}$ is tight on $C([0,T];\mathbb{R}^{d}),$ which is assumed in [12]. But we do not use tightness throughout this paper.
4 Averaging Principle
In this section, we explore the averaging principle for system (2.3). To be precise, we establish the following proposition.
Proposition 4.1.
Let f be the form of either
$f(x,y)=x_{i}y_{k}y_{l}g(x),\quad x=(x_{1},...,x_{d})\in\mathbb{R}^{d},\kern 5.0pty=(y_{1},...,y_{d})\in\mathbb{R}^{d},$ (4.1)
or
$f(x,y)=y_{k}y_{l}g(x),\quad x=(x_{1},...,x_{d})\in\mathbb{R}^{d},\kern 5.0pty=(y_{1},...,y_{d})\in\mathbb{R}^{d},$ (4.2)
where $g\in C^{2}(\mathbb{R}^{d};\mathbb{R})$ is Lipschitz. Under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$), for each $1\leq p<\infty,$
$\lim\limits_{\epsilon\to 0}\mathbb{E}\Big|\int_{0}^{t}f(X^{\epsilon}_{s},Y^{\epsilon}_{s})-\bar{f}(X^{\epsilon}_{s})ds\Big|^{p}=0,$
where $\bar{f}(x):=\int f(x,y)\nu^{x}(dy)$ and $\nu^{x}$ is the unique invariant measure of the following system:
$dY^{x}_{t}=-M(x)Y^{x}_{t}dt+dW_{t}.$ (4.3)
Remark 4.2.
One sees that the non-linearity $F$ vanishes in (4.3). Intuitively speaking, the reason is that, the fast system in (2.3) with frozen slow variable
$dY^{\epsilon,x}_{t}=-\frac{1}{\epsilon^{2}}M(x)Y^{\epsilon,x}_{t}dt+\frac{1}{\epsilon}F(x)dt+\frac{1}{\epsilon}dW_{t},$
becomes
$d\tilde{Y}^{\epsilon,x}_{t}=-M(x)\tilde{Y}^{\epsilon,x}_{t}dt+\epsilon F(x)dt+d\tilde{W}_{t}$ (4.4)
after change-of-variable $\tilde{Y}^{\epsilon,x}_{t}:=Y^{\epsilon,x}_{\epsilon^{2}t},$ where $\tilde{W}_{t}:=\frac{1}{\epsilon}W_{\epsilon^{2}t}\overset{d}{=}W_{t},$ and the term $\epsilon F(x)$ in (4.4) disappears as $\epsilon$ goes to 0.
Remark 4.3.
It is known that [24, Proposition 3.5], under assumption ($\mathbf{A1}$), equation (4.3) admits a unique invariant measure which is Gaussian: $\nu^{x}=\mathcal{N}(0,J(x))$, where $J$ is the solution of (1.3). Consequently,
$\bar{f}(x)=x_{i}g(x)J_{kl}(x)$ (4.5)
if $f$ is of the form (4.1) and
$\bar{f}(x)=g(x)J_{kl}(x)$ (4.6)
if $f$ is of the form (4.2).
Remark 4.4.
It is known that [1, Page 179], under assumption ($\mathbf{A1}$), the Lyapunov equation
$M(x)A(x)+A(x)M^{\top}(x)=B(x)$
admits a unique solution
$A(x)=-\int_{0}^{\infty}e^{-M^{\top}(x)t}B(x)e^{-M(x)t}dt.$
This in particular implies that
$|A(x)|\leq\int_{0}^{\infty}|e^{-M^{\top}(x)t}|\cdot|B(x)|\cdot|e^{-M(x)t}|dt\leq\int_{0}^{\infty}e^{-2\lambda t}|B(x)|dt=\frac{1}{2\lambda}|B(x)|.$
Therefore, if $B$ is of polynomial growth, then so is $A$.
Proof of Proposition 4.1:
We only prove this proposition for $f$ of the form (4.1), since the other case is similar and easier. For each $x\in\mathbb{R}^{d}$, denote the generator of SDE (4.3) by $\mathcal{L}^{x},$ so that
$\mathcal{L}^{x}\phi(x,y)=(D_{y}\phi(x,y),-M(x)y)+\frac{1}{2}\mathrm{Tr}(D_{yy}\phi(x,y)),\quad\phi\in C^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d};\mathbb{R}).$ (4.7)
Consider Poisson equation
$\mathcal{L}^{x}\phi(x,y)=f(x,y)-\bar{f}(x).$ (4.8)
From (4.5) it is straightforward to check that
$\phi(x,y)=x_{i}g(x)(y^{\top}A(x)y-\mathrm{Tr}[A(x)J(x)]),$ (4.9)
where $J$ is the solution of (1.3) and $A$ is the solution of the following Lyapunov equation
$M^{\top}(x)A(x)+A(x)M(x)=\frac{1}{2}(e_{k}e_{l}^{\top}+e_{l}e_{k}^{\top}).$ (4.10)
Here, $(e_{n})_{1\leq n\leq d}$ is the orthonormal basis of $\mathbb{R}^{d}.$ As a consequence of ($\mathbf{A1}$)–($\mathbf{A3}$), $\phi\in C^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d};\mathbb{R}).$ By Itô formula and (4.7),
$d\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t})$
$=$ $\Big(D_{x}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t}),\frac{1}{\epsilon}Y^{\epsilon}_{t}\Big)dt+\Big(D_{y}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t}),-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})Y^{\epsilon}_{t}+\frac{1}{\epsilon}F(X^{\epsilon}_{t})\Big)dt$
$+\frac{1}{2}\mathrm{Tr}\Big[\frac{1}{\epsilon^{2}}D^{2}_{yy}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t})\Big]dt+\Big(\frac{1}{\epsilon}D_{y}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t}),dW_{t}\Big)$
$=$ $\frac{1}{\epsilon}(D_{x}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t}),Y^{\epsilon}_{t})dt+\frac{1}{\epsilon}(D_{y}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t}),F(X^{\epsilon}_{t}))dt$
$+\frac{1}{\epsilon}(D_{y}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t}),dW_{t})+\frac{1}{\epsilon^{2}}\mathcal{L}^{X^{\epsilon}_{t}}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t})dt.$
Integrating on $[0,t]$, rearranging and using (4.8),
$\int_{0}^{t}f(X^{\epsilon}_{s},Y^{\epsilon}_{s})-\bar{f}(X^{\epsilon}_{s})ds$ (4.12)
$=$ $-\int_{0}^{t}\mathcal{L}^{X^{\epsilon}_{s}}\phi(X^{\epsilon}_{s},Y^{\epsilon}_{s})ds$
$=$ $-\epsilon^{2}[\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t})-\phi(0,0)]+\epsilon\int_{0}^{t}(D_{x}\phi(X^{\epsilon}_{s},Y^{\epsilon}_{s}),Y^{\epsilon}_{s})ds$
$+\epsilon\int_{0}^{t}(D_{y}\phi(X^{\epsilon}_{s},Y^{\epsilon}_{s}),F(X^{\epsilon}_{s}))ds+\epsilon\int_{0}^{t}(D_{y}\phi(X^{\epsilon}_{s},Y^{\epsilon}_{s}),dW_{s}).$
With the aid of Proposition 3.1, the proof is finished by checking that $\phi$, $\phi_{x}$ and $\phi_{y}$ are of at most polynomial growth, that is, there exist constants $C>0$ and $q\geq 1$, such that for all $x,y\in\mathbb{R}^{d}$,
$\max\{|\phi(x,y)|,|\phi_{x}(x,y)|,|\phi_{y}(x,y)|\}\leq C(1+|x|^{q}+|y|^{q}).$
Differentiating on both sides of (4.9),
$\phi_{y}(x,y)=2x_{i}g(x)A(x)y,$ (4.13)
and
$\phi_{x_{j}}(x,y)$ (4.14)
$=$ $[\delta_{ij}g(x)+x_{i}\partial_{x_{j}}g(x)]\times[y^{\top}A(x)y-\mathrm{Tr}(A(x)J(x))]$
$+x_{i}g(x)\Big\{y^{\top}[\partial_{x_{j}}A(x)]y-\mathrm{Tr}[(\partial_{x_{j}}A(x))J(x)+A(x)\partial_{x_{j}}J(x)]\Big\}.$
Differentiating on both sides of (1.3) and (4.10), one finds that $\partial_{x_{j}}J(x)$ and $\partial_{x_{j}}A(x)$ satisfy Lyapunov equations
$M(x)\partial_{x_{j}}J(x)+\partial_{x_{j}}J(x)M^{\top}(x)=-([\partial_{x_{j}}M(x)]J(x)+J(x)[\partial_{x_{j}}M^{\top}(x)]).$ (4.15)
and
$M^{\top}(x)\partial_{x_{j}}A(x)+\partial_{x_{j}}A(x)M(x)=-([\partial_{x_{j}}M^{\top}(x)]A(x)+A(x)[\partial_{x_{j}}M(x)]),$ (4.16)
so the polynomial growth property is obtained by combining (4.9), (4.13), (4.14) and Remark 4.4.
5 Proof of the Main Theorem
With preparation above, we prove the main theorem now.
Proof of Theorem 2.2:
By Corollary 3.4 and [9, Theorem A.15], it suffices to prove the pointwise convergence, that is,
$X^{\epsilon}_{t}\to X_{t},\quad\text{in}\kern 5.0ptL^{p}(\Omega,\mathcal{F},P)$
and
$\mathbb{X}^{\epsilon}_{0,t}\to\mathbb{X}_{0,t},\quad\text{in}\kern 5.0ptL^{p}(\Omega,\mathcal{F},P)$
for each $0\leq t\leq T.$
Indeed, from (2.3),
$\epsilon dY^{\epsilon}_{t}=-\frac{1}{\epsilon}M(X^{\epsilon}_{t})Y^{\epsilon}_{t}dt+F(X^{\epsilon}_{t})dt+dW_{t}=-M(X^{\epsilon}_{t})dX^{\epsilon}_{t}+F(X^{\epsilon}_{t})dt+dW_{t}.$
Rearranging,
$dX^{\epsilon}_{t}=M^{-1}(X^{\epsilon}_{t})dW_{t}+M^{-1}(X^{\epsilon}_{t})F(X^{\epsilon}_{t})dt-\epsilon M^{-1}(X^{\epsilon}_{t})dY^{\epsilon}_{t},$ (5.1)
from which
$X^{\epsilon}_{t}$ $=$ $\Big[\int_{0}^{t}M^{-1}(X^{\epsilon}_{s})dW_{s}+\int_{0}^{t}M^{-1}(X^{\epsilon}_{s})F(X^{\epsilon}_{s})ds\Big]-\epsilon\int_{0}^{t}M^{-1}(X^{\epsilon}_{s})dY^{\epsilon}_{s}$ (5.2)
$=:$ $\mathbf{I}(\epsilon,t)-\mathbf{J}(\epsilon,t).$
Plainly,
$\mathbf{I}^{i}(\epsilon,t)=\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})F_{k}(X^{\epsilon}_{s})ds.$ (5.3)
By integration-by-part formula,
$\mathbf{J}^{i}(\epsilon,t)$ (5.4)
$=$ $\epsilon\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})dY^{\epsilon,k}_{s}$
$=$ $\epsilon\sum_{k}[M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\int_{0}^{t}Y^{\epsilon,k}_{s}dM^{-1}_{ik}(X^{\epsilon}_{s})]$
$=$ $\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\epsilon\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})dX^{\epsilon,l}_{s}$
$=$ $\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})ds.$
Combining (5.2), (5.3) and (5.4),
$X^{\epsilon,i}_{t}$
$=$ $\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})F_{k}(X^{\epsilon}_{s})ds$
$+\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})ds-\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}.$
From (1.2),
$X^{i}_{t}=\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})F_{k}(X_{s})ds+\sum_{k,l}\int_{0}^{t}J_{kl}(X_{s})\partial_{x_{l}}M^{-1}_{ik}(X_{s})ds,$
so
$X^{\epsilon,i}_{t}-X^{i}_{t}$ (5.5)
$=$ $\sum_{k}\int_{0}^{t}[M^{-1}_{ik}(X^{\epsilon}_{s})-M^{-1}_{ik}(X_{s})]dW^{k}_{s}+\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})F_{k}(X^{\epsilon}_{s})-M^{-1}_{ik}(X_{s})F_{k}(X_{s})ds$
$-\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}+\sum_{k,l}\int_{0}^{t}[Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})-J_{kl}(X^{\epsilon}_{s})\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})]ds$
$+\sum_{k,l}\int_{0}^{t}[J_{kl}(X^{\epsilon}_{s})\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})-J_{kl}(X_{s})\partial_{x_{l}}M^{-1}_{ik}(X_{s})]ds$
$=:$ $\mathbf{K}_{1}(\epsilon,t)+\mathbf{K}_{2}(\epsilon,t)+\mathbf{K}_{3}(\epsilon,t)+\mathbf{K}_{4}(\epsilon,t)+\mathbf{K}_{5}(\epsilon,t).$
By Burkholder–Davis–Gundy inequality, ($\mathbf{A3}$) and Hölder inequality
$\mathbb{E}|\mathbf{K}_{1}(\epsilon,t)|^{p}$ (5.6)
$\lesssim$ $\sum_{k}\mathbb{E}\Big|\int_{0}^{t}[M^{-1}_{ik}(X^{\epsilon}_{s})-M^{-1}_{ik}(X_{s})]dW^{k}_{s}\Big|^{p}$
$\lesssim$ $\sum_{k}\mathbb{E}\Big(\int_{0}^{t}|M^{-1}_{ik}(X^{\epsilon}_{s})-M^{-1}_{ik}(X_{s})|^{2}ds\Big)^{p/2}$
$\lesssim$ $\sum_{k}\mathbb{E}\Big(\int_{0}^{t}|X^{\epsilon}_{s}-X_{s}|^{2}ds\Big)^{p/2}$
$\lesssim$ $d\mathbb{E}\int_{0}^{t}|X^{\epsilon}_{s}-X_{s}|^{p}ds$
$\lesssim$ $\int_{0}^{t}\mathbb{E}|X^{\epsilon}_{s}-X_{s}|^{p}ds.$
($\mathbf{A3}$) and ($\mathbf{A4}$) imply that
$|\mathbf{K}_{2}(\epsilon,t)|$
$\leq$ $\sum_{k}\int_{0}^{t}|M^{-1}_{ik}(X^{\epsilon}_{s})|\cdot|F_{k}(X^{\epsilon}_{s})-F_{k}(X_{s})|+|F_{k}(X_{s})|\cdot|M^{-1}_{ik}(X^{\epsilon}_{s})-M^{-1}_{ik}(X_{s})|ds$
$\lesssim$ $\int_{0}^{t}|X^{\epsilon}_{s}-X_{s}|ds.$
By Hölder inequality,
$\mathbb{E}|\mathbf{K}_{2}(\epsilon,t)|^{p}\lesssim\mathbb{E}\int_{0}^{t}|X^{\epsilon}_{s}-X_{s}|^{p}ds=\int_{0}^{t}\mathbb{E}|X^{\epsilon}_{s}-X_{s}|^{p}ds.$ (5.7)
By the same argument,
$\mathbb{E}|\mathbf{K}_{5}(\epsilon,t)|^{p}\lesssim\int_{0}^{t}\mathbb{E}|X^{\epsilon}_{s}-X_{s}|^{p}ds.$ (5.8)
Utilizing Proposition 3.1 and ($\mathbf{A3}$),
$\mathbb{E}|\mathbf{K}_{3}(\epsilon,t)|^{p}\lesssim\epsilon^{p}\sum_{k}\sup\limits_{0\leq t\leq T}\mathbb{E}|Y^{\epsilon,k}_{t}|^{p}\lesssim\epsilon^{p}.$ (5.9)
Lastly, from Proposition 4.1 and Remark 4.3, there exists a function $O(\epsilon,t)$ such that
$\mathbb{E}|\mathbf{K}_{4}(\epsilon,t)|^{p}\lesssim O(\epsilon,t),$ (5.10)
and that for each $t\in[0,T],$
$\lim\limits_{\epsilon\to 0}O(\epsilon,t)=0.$
Combining (5.5)–(5.10), we see that
$\mathbb{E}|X^{\epsilon,i}_{t}-X^{i}_{t}|^{p}\lesssim\int_{0}^{t}\mathbb{E}|X^{\epsilon}_{s}-X_{s}|^{p}ds+\epsilon^{p}+O(\epsilon,t).$
Summing up with respect to $i$,
$\mathbb{E}|X^{\epsilon}_{t}-X_{t}|^{p}\lesssim\int_{0}^{t}\mathbb{E}|X^{\epsilon}_{s}-X_{s}|^{p}ds+\epsilon^{p}+O(\epsilon,t).$
By Gronwall inequality,
$\lim\limits_{\epsilon\to 0}\mathbb{E}|X^{\epsilon}_{t}-X_{t}|^{p}=0.$ (5.11)
This shows the convergence of $X^{\epsilon}$. By Proposition 3.1 and dominated convergence theorem,
$\lim\limits_{\epsilon\to 0}\mathbb{E}\Big|\int_{0}^{t}\bar{f}(X^{\epsilon}_{s})-\bar{f}(X_{s})ds\Big|^{p}=0.$ (5.12)
In conjunction with Proposition 4.1,
$\lim\limits_{\epsilon\to 0}\mathbb{E}\Big|\int_{0}^{t}f(X^{\epsilon}_{s},Y^{\epsilon}_{s})-\bar{f}(X_{s})ds\Big|^{p}=0.$ (5.13)
Next, we turn to show the convergence of $\mathbb{X}^{\epsilon}.$ By (5.1),
$\int_{0}^{t}X^{\epsilon}_{s}\otimes dX^{\epsilon}_{s}$ (5.14)
$=$ $\Big(\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})dW_{s}]+\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})F(X^{\epsilon}_{s})]ds\Big)-\epsilon\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})dY^{\epsilon}_{s}]$
$=:$ $\mathbf{A}(\epsilon,t)-\mathbf{B}(\epsilon,t).$
We explore the limit of (5.14) as $\epsilon\to 0$ componentwise. For $\mathbf{A}(\epsilon,t),$ by (5.11), Burkholder–Davis–Gundy inequality and dominated convergence theorem with the help of Proposition 3.1,
$\mathbf{A}^{ij}(\epsilon,t)$ (5.15)
$=$ $\Big(\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})dW_{s}]\Big)_{ij}+\Big(\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})F(X^{\epsilon}_{s})]ds\Big)_{ij}$
$=$ $\int_{0}^{t}X^{\epsilon,i}_{s}\times[M^{-1}(X^{\epsilon}_{s})dW_{s}]^{j}+\int_{0}^{t}X^{\epsilon,i}_{s}\times[M^{-1}(X^{\epsilon}_{s})F(X^{\epsilon}_{s})ds]^{j}$
$=$ $\sum_{k}\int_{0}^{t}X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})F_{k}(X^{\epsilon}_{s})ds$
$\to$ $\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})F_{k}(X_{s})ds,$
where the limit is in $L^{p}(\Omega,\mathcal{F},P).$ For $\mathbf{B}(\epsilon,t),$ integrating by parts,
$\mathbf{B}^{ij}(\epsilon,t)$ (5.16)
$=$ $\epsilon\Big(\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})dY^{\epsilon}_{s}]\Big)_{ij}$
$=$ $\epsilon\int_{0}^{t}X^{\epsilon,i}_{s}[M^{-1}(X^{\epsilon}_{s})dY^{\epsilon}_{s}]^{j}$
$=$ $\epsilon\sum_{k}\int_{0}^{t}X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})dY^{\epsilon,k}_{s}$
$=$ $\epsilon\sum_{k}\Big(X^{\epsilon,i}_{t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\int_{0}^{t}Y^{\epsilon,k}_{s}d[X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})]\Big)$
$=:$ $\mathbf{B}^{ij}_{1}(\epsilon,t)-\mathbf{B}^{ij}_{2}(\epsilon,t).$
By Proposition 3.1 and ($\mathbf{A3}$), $(X^{\epsilon,i}_{t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t})_{0<\epsilon\leq 1}$ is bounded in $L^{p}(\Omega,\mathcal{F},P)$ for each $t\in[0,T],$ so that
$\mathbf{B}_{1}^{ij}(\epsilon,t)\to 0\quad\text{in}\quad L^{p}(\Omega,\mathcal{F},P).$ (5.17)
On the other hand, notice that
$\mathbf{B}_{2}^{ij}(\epsilon,t)$ (5.18)
$=$ $\epsilon\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}d[X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})]$
$=$ $\epsilon\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})dX^{\epsilon,i}_{s}+\epsilon\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}X^{\epsilon,i}_{s}\sum_{l}[\partial_{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{s})dX^{\epsilon,l}_{s}]$
$=$ $\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})ds+\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}X^{\epsilon,i}_{s}\partial_{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{s})ds$
$=:$ $\mathbf{B}_{21}^{ij}(\epsilon,t)+\mathbf{B}_{22}^{ij}(\epsilon,t).$
By (5.13),
$\mathbf{B}_{21}^{ij}(\epsilon,t)\to\sum_{k}\int_{0}^{t}M^{-1}_{jk}(X_{s})J_{ik}(X_{s})ds,$ (5.19)
and
$\mathbf{B}_{22}^{ij}(\epsilon,t)\to\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds,$ (5.20)
where the two limits above are both in $L^{p}(\Omega,\mathcal{F},P).$ Combining (5.15), (5.17), (5.19) and (5.20), we conclude that
$\Big(\int_{0}^{t}X^{\epsilon}_{s}\otimes dX^{\epsilon}_{s}\Big)_{ij}$ (5.21)
$\to$ $\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})F_{k}(X_{s})ds$
$+\sum_{k}\int_{0}^{t}M^{-1}_{jk}(X_{s})J_{ik}(X_{s})ds+\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds,$
in $L^{p}(\Omega,\mathcal{F},P).$
The proof is finished after verifying that the right hand side of (5.21) coincides with $\mathbb{X}^{ij}_{0,t},$ where $\mathbb{X}$ is defined via (2.4). Let us calculate $\mathbb{X}^{ij}_{0,t}.$ Writing (1.2) componentwise, we see that
$\int_{0}^{t}X^{i}_{s}dX^{j}_{s}$
$=$ $\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{ik}(X_{s})F_{k}(X_{s})ds+\sum_{k}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s},$
and
$[X^{i},X^{j}]_{t}=\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds.$
However, multipling both sides of (1.3) by $M^{-1}$ on the left and $(M^{-1})^{\top}$ on the right and then writing it componentwise, we see that
$\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds=\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})J_{kj}(X_{s})+J_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds.$
Combining three equalities above,
$\int_{0}^{t}X^{i}_{s}\circ dX^{j}_{s}$
$=$ $\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{ik}(X_{s})F_{k}(X_{s})ds$
$+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s}+\frac{1}{2}\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})J_{kj}(X_{s})+J_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds.$
Plainly,
$\frac{1}{2}\Big(\int_{0}^{t}J(X_{s})(M^{-1})^{\top}(X_{s})-M^{-1}(X_{s})J(X_{s})ds\Big)_{ij}$
$=$ $\frac{1}{2}\sum_{k}\int_{0}^{t}J_{ik}(X_{s})M^{-1}_{jk}(X_{s})-M^{-1}_{ik}(X_{s})J_{kj}(X_{s})ds.$
Therefore,
$\Big(\int_{0}^{t}X_{s}\otimes\circ dX_{s}\Big)_{ij}+\frac{1}{2}\Big(\int_{0}^{t}J(X_{s})(M^{-1})^{\top}(X_{s})-M^{-1}(X_{s})J(X_{s})ds\Big)_{ij}$ (5.22)
$=$ $\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{ik}(X_{s})F_{k}(X_{s})ds$
$+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}J_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds,$
which coincides with (5.21).

### v2

@@
 
-Abstract    Smoluchowski-Kramers approximation in rough path topology with state-dependent damping is explored.
-The second-order Langevin equation has a form of fast-slow system after suitable change-of-variable, and then its solution is lifted as a rough path in a natural manner.
-Moment estimates of both the original path and the lift are given, followed by which, averaging technique and convergence theorem in rough path topology are used to pass the limit.
-Keywords    Smoluchowski–Kramers approximation, Averaging, Singular pertubation, Rough path theory
+Abstract    We explore the limit of stochastic differential equations driven by some random processes satisfying singularly perturbed second order stochastic differential equations.
+The main tool we employ is the universal limit theorem in rough path theory.
+To this end, we lift the random process as a rough path in a natural manner.
+After suitable change-of-variable, the random process has a form of slow-fast system.
+Moment estimates of both the random process and its lift are given, followed by which, averaging technique and convergence theorem in rough path topology are used to identify the limit.
+Keywords    Smoluchowski–Kramers approximation, Averaging, Singular perturbation, Rough path theory
 2020 Mathematics Subject Classification   60L90, 60H10
 1 Introduction
-Consider the state-dependent second-order Langevin equation
-$\epsilon^{2}\ddot{X}^{\epsilon}_{t}+M(X^{\epsilon}_{t})\dot{X}^{\epsilon}_{t}=F(X^{\epsilon}_{t})+\dot{W}_{t},\kern 5.0ptX^{\epsilon}_{0}=0,\kern 5.0pt\dot{X}^{\epsilon}_{0}=0,$ (1.1)
-defined on a complete stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{0\leq t\leq T},P),$ where $0<T<\infty$ is fixed, $M$ is a matrix-valued function and $F$ is a non-linearity, detailed in Section 2.
-We are interested in the limit of $X^{\epsilon}$ as $\epsilon\to 0,$ which is called Smoluchowski–Kramers (SK) approximation.
-There is plenty of research for the case $M$ is a constant [2, 3, 33, 16, 30, 7], and in this case, it is not surprising that $X^{\epsilon}\to X$ in some sense with
-$M\dot{X}_{t}=F(X_{t})+\dot{W}_{t},\kern 5.0ptX_{0}=0,$
-which can be obtained formally by taking $\epsilon=0$ in (1.1).
-However, when $M$ is not a constant, that is, it is state-dependent, the limit cannot be obtained by simply dropping the small mass term.
-In fact, under certain assumptions, Hottovy et al.[12] showed that
+Let us consider approximation of stochastic differential equations (SDE) of the form
+$dZ^{\epsilon}_{t}=b(Z^{\epsilon}_{t})dt+\sigma(Z^{\epsilon}_{t})dX^{\epsilon}_{t},$ (1.1)
+where $X^{\epsilon}$ is some random process converging to $X$ in some sense. At first glance, one might expect that $Z^{\epsilon}\to Z$ with
+$dZ_{t}=b(Z_{t})dt+\sigma(Z_{t})dX_{t}.$
+However, this is not true in general, as seen in the classical Wong–Zakai approximation, and there is an additional term called Itô–Stratonovich correction, appearing in the limit equation[26].
+In the pioneering work [7], they assumed that the process $X^{\epsilon}$ is governed via the following Langevin equation
+$\epsilon^{2}\ddot{X}^{\epsilon}_{t}+M\dot{X}^{\epsilon}_{t}=\dot{W}_{t},\quad X^{\epsilon}_{0}=0,\kern 5.0pt\dot{X}^{\epsilon}_{0}=0.$ (1.2)
+The process $X^{\epsilon}$ is also called physical Brownian motion (pBm).
+This equation describes the motion of a particle with mass $\epsilon^{2}$ under a random external force $W$, classically assumed to be a standard (mathematical) Brownian motion, and $M$ is the coefficient of friction.
+They firstly showed that $X^{\epsilon}\to X:=M^{-1}W$, which can be formally obtained by setting $\epsilon=0$ in (1.2), and this type of convergence is known as Smoluchowski–Kramers (SK) approximation[3, 4, 16, 30, 7].
+Then, they found that, the solution $Z^{\epsilon}$ of (1.1), converges to a non-trivial limit.
+The reason lies in the fact that, in the context of rough path theory, the Itô lift of $X^{\epsilon}$ does not converge to the Itô lift of $X$.
+However, in the physical model, it commonly happens that the friction coefficient $M$ is not a constant, and there is a non-linear potential term $F$.
+In this paper, we explore the limit of (1.1) with the driving signal $X^{\epsilon}$ modeled by a more general stochastic differential equation
+$\epsilon^{2}\ddot{X}^{\epsilon}_{t}+M(X^{\epsilon}_{t})\dot{X}^{\epsilon}_{t}=F(X^{\epsilon}_{t})+\dot{W}_{t},\kern 5.0ptX^{\epsilon}_{0}=0,\kern 5.0pt\dot{X}^{\epsilon}_{0}=0.$ (1.3)
+We call it a general physical Brownian motion.
+Setting the velocity $V^{\epsilon}_{t}=\dot{X}^{\epsilon}_{t}$ and momentum $P^{\epsilon}_{t}=\epsilon^{2}V^{\epsilon}_{t}$ respectively, (1.3) is equivalent to
+$\begin{cases}dX^{\epsilon}_{t}=V^{\epsilon}_{t}dt,\kern 5.0ptX^{\epsilon}_{0}=0,\\ dV^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})V^{\epsilon}_{t}dt+\frac{1}{\epsilon^{2}}F(X^{\epsilon}_{t})dt+\frac{1}{\epsilon^{2}}dW_{t},\kern 5.0ptV^{\epsilon}_{0}=0,\end{cases}$ (1.4)
+and
+$\begin{cases}dX^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}P^{\epsilon}_{t}dt,\kern 5.0ptX^{\epsilon}_{0}=0,\\ dP^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})P^{\epsilon}_{t}dt+F(X^{\epsilon}_{t})dt+dW_{t},\kern 5.0ptP^{\epsilon}_{0}=0.\end{cases}$ (1.5)
+We also introduce $Y^{\epsilon}_{t}:=\frac{1}{\epsilon}P^{\epsilon}_{t}$ [7, Page 7944], so that
+$\begin{cases}dX^{\epsilon}_{t}=\frac{1}{\epsilon}Y^{\epsilon}_{t}dt,\kern 5.0ptX^{\epsilon}_{0}=0,\\ dY^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})Y^{\epsilon}_{t}dt+\frac{1}{\epsilon}F(X^{\epsilon}_{t})dt+\frac{1}{\epsilon}dW_{t},\kern 5.0ptY^{\epsilon}_{0}=0.\end{cases}$ (1.6)
+The system (1.6) has a form of slow-fast system [6], making it possible to find its limit by averaging method.
+We always assume that equations (1.3) and (1.1) are defined on a complete stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{0\leq t\leq T},P)$ with $0<T<\infty$ fixed, and $W$ is a $d$-dimensional standard Brownian motion.
+Hypothesis on other coefficients are detailed in Section 2.
+We also remark that, since $X^{\epsilon}$ is of finite variation which can be seen from (1.4) , (1.1) is in fact a random ODE.
+The universal limit theorem (also known as continuity of Itô–Lyons map) in rough path theory [19] indicates that, in order to identify the limit of (1.1), it is necessary to find the limit of the lift $\mathbf{X}^{\epsilon}=(X^{\epsilon},\mathbb{X}^{\epsilon})$ of $X^{\epsilon}$, detailed in Section 2.
+To this end, we will firstly calculate the limit of $X^{\epsilon}$.
+In contrast to the constant friction coefficient case, we cannot obtain the limit by simply dropping the small mass term $\epsilon^{2}\ddot{X}^{\epsilon}$ in (1.3).
+In fact, Hottovy et al.[12] showed that, under certain assumptions,
 $\lim_{\epsilon\rightarrow 0}\mathbb{E}\sup\limits_{0\leq t\leq T}|X^{\epsilon}_{t}-X_{t}|^{2}=0,$
@@
 and
-$S_{j}(x)=\sum_{k,l}(\partial_{x_{l}}M^{-1})_{jk}(x)J_{kl}(x),$
+$S_{j}(x)=\sum_{k,l}(\partial_{x_{l}}M^{-1})_{jk}(x)J_{kl}(x),$ (1.8)
 with $J$ the solution of Lyapunov equation
-$M(x)J(x)+J(x)M^{\top}(x)=\mathrm{id}.$ (1.3)
-Since then, there are fruitful work on SK approximation with state-dependent damping.
-For example, Wang et al.[27] explore the limit by averaging on the Fokker-Planck equation associated with the Langevin equation.
-Cerrai and Xi[4] consider the infinite dimensional case.
-For more research in SK approximation in state dependent case, we refer to [5, 11, 17, 28].
-SK approximation is naturally linked to stochastic averaging, initiated by Khaminskii [13].
+$M(x)J(x)+J(x)M^{\top}(x)=\mathrm{I}_{d}.$ (1.9)
+Here $\mathrm{I}_{d}$ is the identity matrix in $\mathbb{R}^{d\times d}.$ For more research in SK approximation in position-dependent friction case, we refer to [5, 11, 17, 28, 27].
+As we see in (1.6), SK approximation is naturally linked to stochastic averaging, initiated by Khaminskii [13].
 Numerous work concerning about stochastic averaging appeared in recent years, see e.g.
@@
 There is also research about averaging principle in rough path framework, see for example in [15, 25].
-In fact, by suitable change-of-variable, (1.1) is transformed to (2.3), which has a form of fast-slow system [6], making it possible to find its limit by averaging method.
-A Poisson equation technique[21, 22, 23] is employed to show the convergence.
-Rough path theory is initially proposed by Lyons[19].
-Although a deterministic theory, it is proved a powerful tool in stochastic analysis.
-See monographs [18, 20, 9, 8] for systematic display of rough path theory.
-In this paper, we aim to find the small mass limit in rough path topology.
-Friz et al [7] identify the limit in rough path topology in the case that $M$ is a constant matrix and $F\equiv 0$.
-They found that, the limit of the "level-2" rough path, converges to a non-trivial limit.
-In particular, when the dimension is $1$, the Itô lift of the original path, converges to the Stratonovich lift of the limit path, which is reminiscent of Wong–Zakai approximation [26].
-They also gave a conjecture for the case when $M$ is not constant, without detailed proof.
-In this paper, the system with both state-dependent friction and non-linearity is explored.
-We identify the limit in the rough path topology, and give rigorous proof.
-This also generalizes the result in [12], in the sense that in our main theorem, the "level-1" path converges in $p$-th moment of $\alpha$-Hölder norm for all $1\leq p<\infty$.
+Therefore, unlike the existing literature, we use a stochastic averaging method to obtain the limit in a much stronger sense.
+The next step is to identify the limit of $\mathbb{X}^{\epsilon}$, which is more delicate.
+After these work, we are able to find the effective approximation of (1.1) by utilizing universal limit theorem.
+Our current article differs from these existing literature in at least three aspects.
+The first is that, our main theorems show the convergence of the lifted process $\mathbf{X}^{\epsilon}$ of $X^{\epsilon}$ in the sense of $p$-th moment of $\alpha$-Hölder inhomogeneous rough path distance with $1/3<\alpha<1/2$.
+This is much stronger than [12] and [27], in which they showed the convergence of the original process $X^{\epsilon}$ in $L^{2}(\Omega;C([0,T];\mathbb{R}^{d}))$ and convergence in distribution, respectively.
+The second is that, in [12], they essentially assume the tightness of $(X^{\epsilon})_{0<\epsilon\leq 1}$ in $C([0,T];\mathbb{R}^{d}).$ In our paper, we drop this assumption, and obtain a much stronger result, that is, $(\mathbf{X}^{\epsilon})_{0<\epsilon\leq 1}$ is tight in $\mathscr{C}^{0,\alpha}([0,T];\mathbb{R}^{d})$, the closure of the smooth rough paths in $\alpha$-Hölder rough path space $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$, which is a corollary of the boundedness of $(\mathbf{X}^{\epsilon})_{0<\epsilon\leq 1}$ in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$.
+In [7], they showed the boundedness by utilizing the Gaussian property of $V^{\epsilon}$ in the constant friction case.
+However, in our setting, $V^{\epsilon}$ is non-Gaussian due to the fact that the friction $M$ depends on the position, which makes the proof more challenging.
+The third is that, as a consequence of the variable friction again, $V^{\epsilon}$ is non-Markovian, so we cannot pass the limit by a direct application of ergodic theorem when calculating the limit of $\mathbb{X}^{\epsilon}$ as in [7].
+To fix this, we prove a strong averaging principle for a specific slow-fast system.
+We will see in Section 4 that our slow system is only locally Lipschitz, which does not satisfy the commonly assumed condition in the aforementioned reference.
+Thanks to the structure of our slow system, we employ Poisson equation technique[21, 22, 23] to show the convergence in $L^{p}$ sense.
 The rest of this paper is organized as follows.
-In Section 2, we recall some basic rough path theory, list some notations, and present our main result.
+In Section 2, we recall some basic rough path theory, list some notations, and present our main results.
 In Section 3, we give moment estimate in rough path norm.
-In Section 4, we display an averaging principle for specific fast-slow system, which is used to pass the limit.
-After these preparation, we prove the main theorem in Section 5.
-2 Preliminary and Main Result
-We recall some basic definitions of rough path theory [8, e.g.]. Let $X:[0,T]\to\mathbb{R}^{d},$ $\mathbb{X}:\Delta_{[0,T]}\to\mathbb{R}^{d\times d}$, where
-$\Delta_{[0,T]}:=\{(s,t)\in[0,T]^{2}:s<t\}.$
+In Section 4, we display an averaging principle for specific slow-fast system, which is used to pass the limit.
+After these preparation, we prove the main theorems in Section 5.
+2 Preliminary and Main Results
+We firstly recall some basic definitions of rough path theory [8, e.g.]. For systematic display of rough path theory, we refer to [18, 20, 9, 8].
+Let $X:[0,T]\to\mathbb{R}^{d},$ $\mathbb{X}:\Delta_{[0,T]}\to\mathbb{R}^{d\times d}$, where
+$\Delta_{[0,T]}:=\{(s,t)\in[0,T]^{2}:s\leq t\}.$
 Given $\frac{1}{3}<\alpha\leq\frac{1}{2},$ a map $\mathbf{X}=(X,\mathbb{X})$ is called an $\alpha$-Hölder rough path over $\mathbb{R}^{d}$, if the following conditions hold:
-(i) $X||_{\alpha}:=\sup_{0\leq s\leq t}\limits\frac{|X_{t}-X_{s}|}{|t-s|^{\alpha}}<\infty$
-(ii) $\mathbb{X}||_{2\alpha}:=\sup_{0\leq s\leq t}\limits\frac{|\mathbb{X}_{s,t}|}{|t-s|^{2\alpha}}<\infty$
-(iii) (Chen relation) $\mathbb{X}_{s,u}=\mathbb{X}_{s,t}+\mathbb{X}_{t,u}+X_{s,t}\otimes X_{t,u},\kern 5.0pt\forall 0\leq s\leq t\leq u\leq T.$
-The set of all $\alpha$-Hölder rough path over $\mathbb{R}^{d}$ is denoted as $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d}).$ Given $\mathbf{X},\mathbf{Y}\in\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d}),$ the $\alpha$-Hölder rough path metric is defined as
+(i) $X||_{\alpha}:=\sup_{0\leq s\leq t\leq T}\limits\frac{|X_{t}-X_{s}|}{|t-s|^{\alpha}}<\infty,$
+(ii) $\mathbb{X}||_{2\alpha}:=\sup_{0\leq s\leq t\leq T}\limits\frac{|\mathbb{X}_{s,t}|}{|t-s|^{2\alpha}}<\infty,$
+(iii) (Chen relation) $\mathbb{X}_{s,u}=\mathbb{X}_{s,t}+\mathbb{X}_{t,u}+X_{s,t}\otimes X_{t,u},\kern 5.0pt0\leq s\leq t\leq u\leq T,$
+where $X_{s,t}:=X_{t}-X_{s}$ for $(s,t)\in\Delta_{[0,T]}.$
+The set of all $\alpha$-Hölder rough path over $\mathbb{R}^{d}$ is denoted by $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d}).$ Note that this space is not a linear space due to the nonlinear Chen relation.
+Given $\mathbf{X},\mathbf{Y}\in\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d}),$ the inhomogeneous $\alpha$-Hölder rough path metric is defined as
 […]
-Let $V^{\epsilon}_{t}=\dot{X}^{\epsilon}_{t}$ and $P^{\epsilon}_{t}=\epsilon^{2}V^{\epsilon}_{t}$ be velocity and momentum, respectively, then (1.1) is equivalent to
-$\begin{cases}dX^{\epsilon}_{t}=V^{\epsilon}_{t}dt,\kern 5.0ptX^{\epsilon}_{0}=0,\\ dV^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})V^{\epsilon}_{t}dt+\frac{1}{\epsilon^{2}}F(X^{\epsilon}_{t})dt+\frac{1}{\epsilon^{2}}dW_{t},\kern 5.0ptV^{\epsilon}_{0}=0,\end{cases}$ (2.1)
+This metric is inhomogeneous with respect to the dilation operator $\delta_{\lambda}$ ($\lambda\in\mathbb{R}$), defined by $\delta_{\lambda}((a,b)):=(\lambda a,\lambda^{2}b),\kern 5.0pta\in\mathbb{R}^{d},b\in\mathbb{R}^{d\times d}.$ Besides, the $\alpha$-Hölder homogeneous rough path norm is defined by
+[…]
+Next we recall concepts of rough differential equations (RDE).
+We adopt the definition in [8, Section 8.7].
+Let $b\in C^{3}_{b}(\mathbb{R}^{e};\mathbb{R}^{e})$, $\sigma\in C^{3}_{b}(\mathbb{R}^{e};\mathbb{R}^{e\times d})$, that is, they are $n$ times continuously differentiable with bounded $k$-th derivative, $k=1,2,3.$ By a solution of the RDE
+$dZ_{t}=b(Z_{t})dt+\sigma(Z_{t})d\mathbf{X}_{t},\quad Z_{0}=\xi,$ (2.1)
+we mean that
+$Z_{t}-Z_{s}=b(Z_{s})(t-s)+\sigma(Z_{s})X_{s,t}+D\sigma(Z_{s})\sigma(Z_{s})\mathbb{X}_{s,t}+R_{s,t},\quad(s,t)\in\Delta_{[0,T]},$
 and
-$\begin{cases}dX^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}P^{\epsilon}_{t}dt,\kern 5.0ptX^{\epsilon}_{0}=0,\\ dP^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})P^{\epsilon}_{t}dt+F(X^{\epsilon}_{t})dt+dW_{t},\kern 5.0ptP^{\epsilon}_{0}=0.\end{cases}$ (2.2)
-Here we also introduce $Y^{\epsilon}_{t}:=\frac{1}{\epsilon}P^{\epsilon}_{t}$ [7, Page 7944], so that
-$\begin{cases}dX^{\epsilon}_{t}=\frac{1}{\epsilon}Y^{\epsilon}_{t}dt,\kern 5.0ptX^{\epsilon}_{0}=0,\\ dY^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})Y^{\epsilon}_{t}dt+\frac{1}{\epsilon}F(X^{\epsilon}_{t})dt+\frac{1}{\epsilon}dW_{t},\kern 5.0ptY^{\epsilon}_{0}=0.\end{cases}$ (2.3)
-For each $0<\epsilon\leq 1,$ define the lift of $X^{\epsilon}$ by
-$\mathbb{X}^{\epsilon}_{s,t}:=\int_{s}^{t}X^{\epsilon}_{s,u}\otimes dX^{\epsilon}_{u},\quad(s,t)\in\Delta_{[0,T]},$
-where the stochastic integral is Itô integral, and set $\mathbf{X}^{\epsilon}=(X^{\epsilon},\mathbb{X}^{\epsilon}).$ It is straightforward to check that $\mathbf{X}^{\epsilon}$ satisfies Chen relation.
-Moreover, by Kolmogorov criterion for rough paths [8, Theorem 3.1] and Proposition 3.3 below, $\mathbf{X}^{\epsilon}$ lies in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$ for all $\frac{1}{3}<\alpha<\frac{1}{2}.$
+$\lim\limits_{|\mathcal{P}|\to 0}\sup\limits_{\mathcal{P}\subset[0,T]}\sum\limits_{[s,t]\in\mathcal{P}}|R_{s,t}|=0,$
+here $\mathcal{P}\subset[0,T]$ means $\mathcal{P}$ is a partition of $[0,T]$, and $|\mathcal{P}|$ is the mesh of $\mathcal{P}$.
+It turns out that, the Itô–Lyons map of (2.1) is locally Lipschitz with respect to $\rho_{\alpha}$[8, Thereom 8.5].
+To be precise, if $\tilde{Z}$ is the solution of
+$d\tilde{Z}_{t}=b(\tilde{Z}_{t})dt+\sigma(\tilde{Z}_{t})d\tilde{\mathbf{X}}_{t},\quad\tilde{Z}_{0}=\tilde{\xi}$, (2.2)
+[…]
+[…]
+With preliminaries above, we are ready to display our main results. For each $0<\epsilon\leq 1,$ define the second-order process by
+$\mathbb{X}^{\epsilon}_{s,t}:=\int_{s}^{t}X^{\epsilon}_{s,u}\otimes dX^{\epsilon}_{u},\quad(s,t)\in\Delta_{[0,T]}.$ (2.3)
+Since $X^{\epsilon}$ is of finite variation, the integral above is in the sense of Riemann–Stieltjes.
+Setting $\mathbf{X}^{\epsilon}=(X^{\epsilon},\mathbb{X}^{\epsilon}),$ it is straightforward to check that $\mathbf{X}^{\epsilon}$ satisfies Chen relation, and that $\mathbf{X}^{\epsilon}$ lies in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$ for all $\frac{1}{3}<\alpha<\frac{1}{2}.$
 We impose the following assumptions:
@@
 Well-posedness of (1.7) and (1.6) are guaranteed by ($\mathbf{A3}$) and ($\mathbf{A4}$).
-Now we state the main result.
+Now we state our main theorems.
 Theorem 2.2.
@@
 Remark 2.3.
-By Kolmogorov criterion for rough paths again, one verifies that $\mathbf{X}$ lies in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$ for all $\frac{1}{3}<\alpha<\frac{1}{2}.$
+By Kolmogorov criterion for rough paths[8, Theorem 3.1], one verifies that $\mathbf{X}$ lies in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$ for all $\frac{1}{3}<\alpha<\frac{1}{2}.$
+An application of continuity of Itô–Lyons map with careful computation, detailed in Section 5, yield the following theorem.
+Theorem 2.4.
+Let $b\in C^{3}_{b}(\mathbb{R}^{e};\mathbb{R}^{e})$ and $\sigma\in C^{3}_{b}(\mathbb{R}^{e};\mathbb{R}^{e\times d}).$ Consider random ODE
+$dZ^{\epsilon}_{t}=b(Z^{\epsilon}_{t})dt+\sigma(Z^{\epsilon}_{t})dX^{\epsilon}_{t},\quad Z^{\epsilon}_{0}=\xi.$
+Under assumptions $(\mathbf{A1})$$(\mathbf{A4})$, $Z^{\epsilon}\to Z$ in $C^{\alpha}([0,T];\mathbb{R}^{e})$ in probability with $Z_{0}=\xi$ and
+$dZ^{i}_{t}$
+$=$ $[b_{i}(Z_{t})+\sum\limits_{j,k,l,k^{\prime}}\partial_{z_{k}}\sigma_{ij}(Z_{t})\sigma_{kl}(Z_{t})J_{lk^{\prime}}(X_{t})M^{-1}_{jk^{\prime}}(X_{t})$
+$+\sum_{j}\sigma_{ij}(Z_{t})S_{j}(X_{t})+\sum_{j,k}\sigma_{ij}(Z_{t})M^{-1}_{jk}(X_{t})F_{k}(X_{t})]dt+\sum_{j,k}\sigma_{ij}(Z_{t})M^{-1}_{jk}(X_{t})dW^{k}_{t},$
+where $S_{j}$ is defined in (1.8).
 We end this section with a collection of frequently used notation.
@@
 $|\cdot|$: Euclidean norm of either a vector or a matrix.
-$A^{\top}$: Transpose of a matrix or a vector $A$.
-$\mathrm{id}$: Identity matrix.
+$A^{\top}$: Transpose of a matrix or a vector AA.
+$\mathrm{I}_{d}$: Identity matrix of $d\times d$.
 $(x,y)$: Inner product of two vectors $x$ and $y$.
 $Df,D^{2}f$: Gradient and Hessian of a function $f:\mathbb{R}^{d}\to\mathbb{R}$.
-$C^{k}(U;V)$: The set consists of $k$-th continuously differentiable functions from $U$ to $V$, where $k\in\mathbb{N},$ and $U$, $V$ are Banach spaces.
+$C^{k}(U;V)$: The set consists of $k$-th continuously differentiable functions from $U$ to $V$, where $k\in\mathbb{N},$ and $U$, $V$ are finite-dimensional Banach spaces.
 $a\lesssim b$: There exists a constant $C>0$ such that $a\leq Cb.$ Throughout the paper, $C$ never depends on $\epsilon$.
@@
 3 Moment Estimates
-In this section, we give several moment estimate for system (2.3). We begin with bounding the supremum norm.
+In this section, we give several moment estimates for system (1.6).
 Proposition 3.1.
 Under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$), for each $1\leq p<\infty,$
-$\sup\limits_{0<\epsilon\leq 1}\sup\limits_{0\leq t\leq T}\mathbb{E}|Y^{\epsilon}_{t}|^{p}<\infty,\kern 5.0pt\sup\limits_{0<\epsilon\leq 1}\mathbb{E}\sup\limits_{0\leq t\leq T}|X^{\epsilon}_{t}|^{p}<\infty.$
+$\sup\limits_{0<\epsilon\leq 1}\sup\limits_{0\leq t\leq T}\mathbb{E}|Y^{\epsilon}_{t}|^{p}<\infty.$
 Proof.
-We firstly show the boundedness for $Y^{\epsilon}$. For each $p\geq 2,$ set $\phi(y)=|y|^{p},$ then
-$D\phi(y)=p|y|^{p-2}y,\quad D^{2}\phi(y)=p(p-2)|y|^{p-4}yy^{\top}+p|y|^{p-2}\mathrm{id}.$
-Apply Itô formula to $\phi(Y^{\epsilon}_{t})$ and use ($\mathbf{A1}$) and ($\mathbf{A4}$) to find that
+For each $p\geq 2,$ set $\phi(y)=|y|^{p},$ then
+$D\phi(y)=p|y|^{p-2}y,\quad D^{2}\phi(y)=p(p-2)|y|^{p-4}yy^{\top}+p|y|^{p-2}\mathrm{I}_{d}.$
+Applying Itô formula to $\phi(Y^{\epsilon}_{t})$,
 $|Y^{\epsilon}_{t}|^{p}$ (3.1)
@@
 $=$ $-\frac{p}{\epsilon^{2}}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-2}(Y^{\epsilon}_{s})^{\top}M(X^{\epsilon}_{s})Y^{\epsilon}_{s}ds+\frac{p}{\epsilon}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-2}(Y^{\epsilon}_{s})^{\top}F(X^{\epsilon}_{s})ds+H^{\epsilon}_{t}$
-$+\frac{1}{2\epsilon^{2}}\int_{0}^{t}\mathrm{Tr}[p(p-2)|Y^{\epsilon}_{s}|^{p-4}Y^{\epsilon}_{s}(Y^{\epsilon}_{s})^{\top}+p|Y^{\epsilon}_{s}|^{p-2}\mathrm{id}]$
-$\lesssim$ $-\frac{1}{\epsilon^{2}}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p}ds+\frac{1}{\epsilon}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-1}ds+H^{\epsilon}_{t}+\frac{1}{\epsilon^{2}}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-2}ds,$
-where $H^{\epsilon}$ is a martingale. Taking expectation and applying Hölder inequality,
-$\mathbb{E}|Y^{\epsilon}_{t}|^{p}\lesssim-\frac{1}{\epsilon^{2}}\int_{0}^{t}\mathbb{E}|Y^{\epsilon}_{s}|^{p}ds+\frac{1}{\epsilon^{2}}\int_{0}^{t}\mathbb{E}|Y^{\epsilon}_{s}|^{p-2}ds.$ (3.2)
+$+\frac{1}{2\epsilon^{2}}\int_{0}^{t}\mathrm{Tr}[p(p-2)|Y^{\epsilon}_{s}|^{p-4}Y^{\epsilon}_{s}(Y^{\epsilon}_{s})^{\top}+p|Y^{\epsilon}_{s}|^{p-2}\mathrm{I}_{d}]ds,$
+where $H^{\epsilon}$ is a martingale. Taking expectation, differentiating and applying $(\mathbf{A1})$ and $(\mathbf{A4})$,
+$\frac{d}{dt}\mathbb{E}|Y^{\epsilon}_{t}|^{p}\lesssim-\frac{1}{\epsilon^{2}}\mathbb{E}|Y^{\epsilon}_{t}|^{p}+\frac{1}{\epsilon^{2}}\mathbb{E}|Y^{\epsilon}_{t}|^{p-2}.$ (3.2)
 First for $p=2$, by the comparison principle,
@@
 for all even $p$, and hence for all $1\leq p<\infty.$
-Next, we show the boundedness for $X^{\epsilon}$. Consider the following random ODE in $\mathbb{R}^{d\times d}$
-$\dot{\phi}^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})\phi^{\epsilon}_{t},\kern 5.0pt\phi^{\epsilon}_{0}=\mathrm{id}.$
-For each $r\in[0,T],$ set $z^{\epsilon,r}_{s}=\phi^{\epsilon}_{s}(\phi^{\epsilon}_{r})^{-1},\kern 5.0pts\in[r,T].$ It is showed in [27, Page 6] that, under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$),
-$X^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{0}^{s}z^{\epsilon,r}_{s}dW_{r}ds+\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{0}^{s}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})drds.$ (3.3)
-and that
-$|z^{\epsilon,r}_{s}|\leq e^{-\epsilon^{-2}\lambda(s-r)},\quad 0\leq r\leq s\leq T.$ (3.4)
-For notation simplicity, we always assume that $\lambda=1$. Then, from (3.4)
-$\Big|\int_{r}^{t}z^{\epsilon,r}_{s}ds\Big|\leq\int_{r}^{t}|z^{\epsilon,r}_{s}|ds\leq\epsilon^{2}.$ (3.5)
-Notice that by stochastic Fubini theorem, we have
-$X^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}dsdW_{r}+\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})dsdr.$ (3.6)
-An application of Burkholder–Davis–Gundy inequality and (3.5) yield
-$\frac{1}{\epsilon^{2p}}\mathbb{E}\sup\limits_{0\leq t\leq T}\Big|\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}dsdW_{r}\Big|^{p}$ (3.7)
-$\lesssim$ $\frac{1}{\epsilon^{2p}}\mathbb{E}\Big(\int_{0}^{T}\Big|\int_{r}^{t}z^{\epsilon,r}_{s}ds\Big|^{2}dr\Big)^{p/2}$
-$\lesssim$ $1.$
-By ($\mathbf{A4}$) and (3.5), for each $t\in[0,T],$
-$\frac{1}{\epsilon^{2p}}\Big|\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})dsdr\Big|^{p}$ (3.8)
-$\leq$ $\frac{1}{\epsilon^{2p}}\Big(\int_{0}^{t}\int_{r}^{t}|z^{\epsilon,r}_{s}|dsdr\Big)^{p}$
-$\leq$ $T^{p}.$
-Taking supremum and expectation,
-$\frac{1}{\epsilon^{2p}}\mathbb{E}\sup\limits_{0\leq t\leq T}\Big|\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})dsdr\Big|^{p}\lesssim 1.$ (3.9)
-The result follows from combining (3.6), (3.7) and (3.9). ∎
-We need the following elementary inequality, which is a special case of Minkowski integral inequality[10, Page 13]. For completeness, we give its proof.
-Lemma 3.2.
-Let $G$ be a random process on $[0,T],$ $a,b]\subset[0,T],$ $q\in\mathbb{N},$ then
-$\Big[\mathbb{E}\Big(\int_{a}^{b}|G(r)|^{2}dr\Big)^{q}\Big]^{1/q}\leq\int_{a}^{b}\Big(\mathbb{E}|G(r)|^{2q}\Big)^{1/q}dr.$
-Proof.
-Applying Hölder inequality,
-$\mathbb{E}\Big(\int_{a}^{b}|G(r)|^{2}dr\Big)^{q}$
-$=$ $\mathbb{E}\int_{[a,b]^{q}}|G(r_{1})|^{2}\cdots|G(r_{q})|^{2}dr_{1}\cdots dr_{q}$
-$=$ $\int_{[a,b]^{q}}\mathbb{E}(|G(r_{1})|^{2}\cdots|G(r_{q})|^{2})dr_{1}\cdots dr_{q}$
-$\leq$ $\int_{[a,b]^{q}}(\mathbb{E}|G(r_{1})|^{2q})^{1/q}\cdots(\mathbb{E}|G(r_{q})|^{2q})^{1/q}dr_{1}\cdots dr_{q}$
-$=$ $\Big(\int_{a}^{b}(\mathbb{E}|G(r)|^{2q})^{1/q}dr\Big)^{q}.$
 The next proposition concerns about Hölder norm of the path $X^{\epsilon}$ and its lift.
@@
 Proof.
-In view of Jensen inequality, it is sufficient to prove the proposition for all even $p.$ We first prove (i). From (3.3) and stochastic Fubini theorem,
+We give estimates by dichotomy. From (1.6),
+$\epsilon dY^{\epsilon}_{t}=-\frac{1}{\epsilon}M(X^{\epsilon}_{t})Y^{\epsilon}_{t}dt+F(X^{\epsilon}_{t})dt+dW_{t}=-M(X^{\epsilon}_{t})dX^{\epsilon}_{t}+F(X^{\epsilon}_{t})dt+dW_{t}.$
+Rearranging,
+$dX^{\epsilon}_{t}=M^{-1}(X^{\epsilon}_{t})dW_{t}+M^{-1}(X^{\epsilon}_{t})F(X^{\epsilon}_{t})dt-\epsilon M^{-1}(X^{\epsilon}_{t})dY^{\epsilon}_{t},$ (3.3)
+Case 1: $\epsilon\geq\sqrt{t-s}$.
+By (1.6), Hölder inequality, Fubini theorem and Proposition 3.1,
+$\mathbb{E}|X^{\epsilon}_{t}-X^{\epsilon}_{s}|^{p}$ $\leq$ $\frac{1}{\epsilon^{p}}\mathbb{E}\Big(\int_{s}^{t}|Y^{\epsilon}_{u}|du\Big)^{p}$ (3.4)
+$\leq$ $\frac{1}{\epsilon^{p}}(t-s)^{p-1}\int_{s}^{t}\mathbb{E}|Y^{\epsilon}_{u}|^{p}du$
+$\leq$ $\frac{1}{\epsilon^{p}}(t-s)^{p}\leq(t-s)^{p/2}.$
+Case 2: $\epsilon<\sqrt{t-s}$.
+From (3.3),
 $X^{\epsilon}_{t}-X^{\epsilon}_{s}$ (3.5)
-$=$ $\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}z^{\epsilon,r}_{u}dW_{r}du+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})drdu$
-$=$ $\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}dudW_{r}+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big]$
-$+\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big].$
-Now we deal with four terms above respectively. By Burkholder–Davis–Gundy inequality,
-$\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}\lesssim\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}.$ (3.11)
-From (3.4) and Hölder inequality
-$\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}$
-$\leq$ $\Big|\int_{s}^{t}e^{-\epsilon^{-2}(u-r)}du\Big|^{2}$
-$\leq$ $(t-s)\int_{s}^{t}e^{-2\epsilon^{-2}(u-r)}du.$
-Therefore,
-$\int_{0}^{s}\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}dr\lesssim(t-s)\int_{0}^{s}\int_{s}^{t}e^{-2\epsilon^{-2}(u-r)}dudr\lesssim\epsilon^{4}(t-s).$ (3.12)
-Substituting the above inequality into (3.11),
-$\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}\lesssim\epsilon^{2p}|t-s|^{p/2}.$ (3.13)
-By Burkholder–Davis–Gundy inequality and (3.4),
-$\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}$ (3.14)
-$\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}z^{\epsilon,r}_{u}du\Big|^{2}\Big)^{p/2}$
-$\leq$ $\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}e^{-\epsilon^{-2}(u-r)}du\Big|^{2}dr\Big)^{p/2}$
-$\leq$ $\Big(\int_{s}^{t}\epsilon^{4}dr\Big)^{p/2}$
-$\leq$ $\epsilon^{2p}(t-s)^{p/2}.$
-On the other hand, by ($\mathbf{A4}$), Hölder inequality and (3.12),
-$\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.15)
-$\lesssim$ $\mathbb{E}\Big(\int_{0}^{s}\int_{s}^{t}|z^{\epsilon,r}_{u}|dudr\Big)^{p}$
-$=$ $\mathbb{E}\Big\{\Big[\Big(\int_{0}^{s}\int_{s}^{t}|z^{\epsilon,r}_{u}|dudr\Big)^{2}\Big]^{p/2}\Big\}$
-$\lesssim$ $\mathbb{E}\Big\{\Big[\int_{0}^{s}\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big]^{p/2}\Big\}$
-$\lesssim$ $\epsilon^{2p}|t-s|^{p/2}.$
-Lastly,
-$\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.16)
-$\lesssim$ $\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}e^{-\epsilon^{-2}(u-r)}dudr\Big|^{p}$
-$\leq$ $\Big|\int_{s}^{t}\epsilon^{2}dr\Big|^{p}$
-$\leq$ $\epsilon^{2p}(t-s)^{p}.$
-Substituting (3.13)–(3.16) into (3.10),
-$\mathbb{E}|X^{\epsilon}_{t}-X^{\epsilon}_{s}|^{p}\lesssim|t-s|^{p/2}.$ (3.17)
-We turn to prove (ii).
-By triangular inequality, it is sufficient to consider the $(i,j)$-entry $\int_{s}^{t}X^{\epsilon,i}_{s,u}dX^{\epsilon,j}_{u}$ of $\int_{s}^{t}X^{\epsilon}_{s,u}\otimes dX^{\epsilon}_{u}.$ But from (3.3),
-$dX^{\epsilon,j}_{t}=\frac{1}{\epsilon^{2}}\sum_{k}\int_{0}^{t}z^{\epsilon,r}_{t}(j,k)dB^{k}_{r}dt+\frac{1}{\epsilon^{2}}\sum_{k}\int_{0}^{t}z^{\epsilon,r}_{t}(j,k)F_{k}(X^{\epsilon}_{r})drdt,$
-where $z^{\epsilon,r}_{t}(j,k)$ is the ($j,k$)-entry of $z^{\epsilon,r}_{t}$, so
-$\int_{s}^{t}X^{\epsilon,i}_{s,u}dX^{\epsilon,j}_{u}=\frac{1}{\epsilon^{2}}\sum_{k}\int_{s}^{t}\int_{0}^{u}X^{\epsilon,i}_{s,u}z^{\epsilon,r}_{u}(j,k)dB^{k}_{r}du+\frac{1}{\epsilon^{2}}\sum_{k}\int_{s}^{t}\int_{0}^{u}X^{\epsilon,i}_{s,u}z^{\epsilon,r}_{u}(j,k)F_{k}(X^{\epsilon}_{r})drdu.$
-Without loss of generality, we assume that the system is one-dimensional, that is, we write
-$\int_{s}^{t}X^{\epsilon}_{s,u}dX^{\epsilon}_{u}=\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dW_{r}du+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})drdu.$ (3.18)
-By stochastic Fubini theorem,
-$\int_{s}^{t}X^{\epsilon}_{s,u}dX^{\epsilon}_{u}$ (3.19)
-$=$ $\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}\Big]$
-$+\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big]$
-$=:$ $I+J.$
-We firstly estimate $I$. Using Burkholder–Davis–Gundy inequality,
-$\mathbb{E}|I|^{p}$ (3.20)
-$\lesssim$ $\frac{1}{\epsilon^{2p}}\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}+\frac{1}{\epsilon^{2p}}\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}$
-$\lesssim$ $\frac{1}{\epsilon^{2p}}\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}+\frac{1}{\epsilon^{2p}}\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}.$
-Applying Lemma 3.2 to $G(r)=\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du$ and $q=p/2,$
-$\Big\{\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}\leq\int_{0}^{s}\Big(\mathbb{E}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{2/p}dr.$ (3.21)
-Using Hölder inequality and taking $p$-th power,
-$\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\leq\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}|^{2}du\Big)^{p/2}\Big(\int_{s}^{t}|z^{\epsilon,r}_{u}|^{2}\Big)^{p/2}.$ (3.22)
-Applying Lemma 3.2 to $G(u)=|X^{\epsilon}_{s,u}|$ and $q=p/2,$ and then using (3.17),
-$\Big\{\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}|^{2}du\Big)^{p/2}\Big\}^{2/p}$
-$\leq$ $\int_{s}^{t}(\mathbb{E}|X^{\epsilon}_{s,u}|^{p})^{2/p}du$
-$\lesssim$ $\int_{s}^{t}|u-s|du$
-$=$ $\frac{(t-s)^{2}}{2}.$
-Taking ($p/2$)-th power,
-$\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}|^{2}du\Big)^{p/2}\lesssim|t-s|^{p}.$ (3.23)
-On the other hand, from (3.4) we have
-$\int_{s}^{t}|z^{\epsilon,r}_{u}|^{2}du\leq\int_{s}^{t}e^{-2\epsilon^{-2}(u-r)}du\leq\frac{\epsilon^{2}}{2}e^{-2\epsilon^{-2}(s-r)}.$
-Taking $(p/2)$-th power,
-$\Big(\int_{s}^{t}|z^{\epsilon,r}_{u}|^{2}du\Big)^{p/2}\lesssim\epsilon^{p}e^{-p\epsilon^{-2}(s-r)}.$ (3.24)
-Combining (3.22), (3.23) and (3.24),
-$\mathbb{E}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\leq\epsilon^{p}e^{-p\epsilon^{-2}(s-r)}|t-s|^{p}.$
-Taking ($2/p$)-th power and substituting into (3.21),
-$\Big\{\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}$
-$\lesssim$ $\int_{0}^{s}\epsilon^{2}e^{-2\epsilon^{-2}(s-r)}|t-s|^{2}dr$
-$=$ $\epsilon^{2}(t-s)^{2}\int_{0}^{s}e^{-2\epsilon^{-2}(s-r)}dr$
-$\leq$ $\epsilon^{4}(t-s)^{2}.$
-Taking ($p/2$)-th power on the above inequality,
-$\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\lesssim\epsilon^{2p}(t-s)^{p}.$ (3.25)
-We deal with the other term in a similar fashion. Applying Lemma 3.2 to $G(r)=\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du$ and $q=p/2,$
-$\Big\{\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}\leq\int_{s}^{t}\Big(\mathbb{E}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{2/p}dr.$ (3.26)
-For fixed $r\in[s,t],$ using Lemma 3.2 to $G(u)=\sqrt{|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|},u\in[r,t]$ and $q=p$,
-$\Big(\mathbb{E}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{1/p}$ (3.27)
-$\leq$ $\Big(\mathbb{E}\Big[\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|du\Big]^{p}\Big)^{1/p}$
-$\leq$ $\int_{r}^{t}\Big(\mathbb{E}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|^{p})^{1/p}du.$
-By (3.4) and (3.17),
-$\mathbb{E}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|^{p}\leq e^{-p\epsilon^{-2}(u-r)}\mathbb{E}|X^{\epsilon}_{s,u}|^{p}\leq e^{-p\epsilon^{-2}(u-r)}|u-s|^{p/2}.$
-Taking $(1/p)$-th power and substituting it into (3.27),
-$\Big(\mathbb{E}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{1/p}\leq\int_{r}^{t}e^{-\epsilon^{-2}(u-r)}|u-s|^{1/2}du\leq\epsilon^{2}|t-s|^{1/2}.$ (3.28)
-Taking square and substituting it into (3.26),
-$\Big\{\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}\leq\int_{s}^{t}\epsilon^{4}|t-s|dr=\epsilon^{4}|t-s|^{2}.$ (3.29)
-Taking $(p/2)$-th power,
-$\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\leq\epsilon^{2p}|t-s|^{p}.$ (3.30)
-Combining (3.20), (3.25) and (3.30),
-$\mathbb{E}|I|^{p}\lesssim|t-s|^{p}.$ (3.31)
-It remains to deal with $J$. Writing
-$\Big|\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.32)
-$\lesssim$ $\Big(\int_{0}^{s}\int_{s}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{p}$
-$=$ $\Big[\Big(\int_{0}^{s}\int_{s}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{2}\Big]^{p/2}$
-$\lesssim$ $\Big[\int_{0}^{s}\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|du\Big)^{2}dr\Big]^{p/2},$
+$=$ $\int_{s}^{t}M^{-1}(X^{\epsilon}_{u})dW_{u}+\int_{s}^{t}M^{-1}(X^{\epsilon}_{u})F(X^{\epsilon}_{u})du-\epsilon\int_{s}^{t}M^{-1}(X^{\epsilon}_{u})dY^{\epsilon}_{u}$
+$=:$ $I^{\epsilon,1}(s,t)+I^{\epsilon,2}(s,t)+I^{\epsilon,3}(s,t).$
+By Burkholder–Davis–Gundy inequality and ($\mathbf{A4}$),
+$\mathbb{E}|I^{\epsilon,1}(s,t)|^{p}\lesssim\mathbb{E}\Big(\int_{s}^{t}|M^{-1}(X^{\epsilon}_{u})|^{2}du\Big)^{p/2}\lesssim(t-s)^{p/2}.$ (3.6)
+$\mathbb{E}|I^{\epsilon,2}(s,t)|^{p}\lesssim\mathbb{E}\Big(\int_{s}^{t}|M^{-1}(X^{\epsilon}_{u})F(X^{\epsilon}_{u})|du\Big)^{p}\lesssim(t-s)^{p}\lesssim(t-s)^{p/2}.$ (3.7)
+Applying integration-by-part formula and (1.6),
+$I^{\epsilon,3}_{i}(s,t)$ (3.8)
+$=$ $\epsilon\sum_{j}\int_{s}^{t}M^{-1}_{ij}(X^{\epsilon}_{u})dY^{\epsilon,j}_{u}$
+$=$ $\epsilon\sum_{j}[M^{-1}_{ij}(X^{\epsilon}_{t})Y^{\epsilon,j}_{t}-M^{-1}_{ij}(X^{\epsilon}_{s})Y^{\epsilon,j}_{s}]-\epsilon\sum_{j,l}\int_{s}^{t}Y^{\epsilon,j}_{u}dM^{-1}_{ij}(X^{\epsilon}_{u})$
+$=$ $\epsilon\sum_{j}[M^{-1}_{ij}(X^{\epsilon}_{t})Y^{\epsilon,j}_{t}-M^{-1}_{ij}(X^{\epsilon}_{s})Y^{\epsilon,j}_{s}]-\epsilon\sum_{j,l}\int_{s}^{t}Y^{\epsilon,j}_{u}\partial_{x_{l}}M^{-1}_{ij}(X^{\epsilon}_{u})dX^{\epsilon,l}_{u}$
+$=$ $\epsilon\sum_{j}[M^{-1}_{ij}(X^{\epsilon}_{t})Y^{\epsilon,j}_{t}-M^{-1}_{ij}(X^{\epsilon}_{s})Y^{\epsilon,j}_{s}]-\sum_{j,l}\int_{s}^{t}Y^{\epsilon,j}_{u}Y^{\epsilon,l}_{u}\partial_{x_{l}}M^{-1}_{ij}(X^{\epsilon}_{u})du$
+$=:$ $J^{\epsilon,1}_{i}(s,t)+J^{\epsilon,2}_{i}(s,t).$
+Since
+$J^{\epsilon,1}_{i}(s,t)=\epsilon\sum_{j}[M^{-1}_{ij}(X^{\epsilon}_{t})(Y^{\epsilon,j}_{t}-Y^{\epsilon,j}_{s})+Y^{\epsilon,j}_{s}(M^{-1}_{ij}(X^{\epsilon}_{t})-M^{-1}_{ij}(X^{\epsilon}_{s}))],$
+an application of ($\mathbf{A3}$), Proposition 3.1 and the dichotomy assumption yields
+$\mathbb{E}|J^{\epsilon,1}_{i}(s,t)|^{p}$ (3.9)
+$\lesssim$ $\epsilon^{p}\mathbb{E}(|M^{-1}_{ij}(X^{\epsilon}_{t})|^{p}|Y^{\epsilon,j}_{t}-Y^{\epsilon,j}_{s}|^{p})+\epsilon^{p}\mathbb{E}(|Y^{\epsilon,j}_{s}|^{p}|M^{-1}_{ij}(X^{\epsilon}_{t})-M^{-1}_{ij}(X^{\epsilon}_{s})|^{p})$
+$\lesssim$ $\epsilon^{p}\leq(t-s)^{p/2}.$
+By Hölder inequality, ($\mathbf{A3}$) and Proposition 3.1,
+$\mathbb{E}|J^{\epsilon,2}_{i}(s,t)|^{p}$ $\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}|Y^{\epsilon,j}_{u}Y^{\epsilon,l}_{u}\partial_{x_{l}}M^{-1}_{ij}(X^{\epsilon}_{u})|du\Big)^{p}$ (3.10)
+$\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}|Y^{\epsilon,j}_{u}Y^{\epsilon,l}_{u}|du\Big)^{p}$
+$\leq$ $(t-s)^{p-1}\int_{s}^{t}\mathbb{E}|Y^{\epsilon,j}_{u}Y^{\epsilon,l}_{u}|^{p}du$
+$\lesssim$ $(t-s)^{p}\lesssim(t-s)^{p/2}.$
+Combining (3.8)–(3.10),
+$\mathbb{E}|I^{\epsilon,3}_{i}(s,t)|^{p}\lesssim(t-s)^{p/2}.$ (3.11)
+Combining (3.5)–(3.7) and (3.11), we have
+$\mathbb{E}|X^{\epsilon}_{t}-X^{\epsilon}_{s}|^{p}\lesssim(t-s)^{p/2},$ (3.12)
+and this finishes the proof of (i). Next we turn to (ii).
+Case 1: $\epsilon\geq\sqrt{t-s}$.
+From (1.6) we see that
+$\mathbb{X}^{\epsilon}_{s,t}=\frac{1}{\epsilon}\int_{s}^{t}X^{\epsilon}_{s,u}\otimes Y^{\epsilon}_{u}du.$
+By Hölder inequality, Proposition 3.1 and Proposition 3.2 (i),
+$\mathbb{E}|\mathbb{X}^{\epsilon}_{s,t}|^{p}$ $\leq$ $\frac{1}{\epsilon^{p}}(t-s)^{p-1}\int_{s}^{t}\mathbb{E}|X^{\epsilon}_{s,u}\otimes Y^{\epsilon}_{u}|^{p}du$ (3.13)
+$\leq$ $\frac{1}{\epsilon^{p}}(t-s)^{p-1}\int_{s}^{t}\sqrt{\mathbb{E}|X^{\epsilon}_{s,u}|^{2p}}\sqrt{\mathbb{E}|Y^{\epsilon}_{u}|^{2p}}du$
+$\lesssim$ $\frac{1}{\epsilon^{p}}(t-s)^{p-1}\int_{s}^{t}(u-s)^{p/2}du$
+$\lesssim$ $\frac{1}{\epsilon^{p}}(t-s)^{3p/2}\leq(t-s)^{p}.$
+Case 2: $\epsilon<\sqrt{t-s}$.
+By (3.3),
+$\mathbb{X}^{\epsilon}_{s,t}$ (3.14)
+$=$ $\int_{s}^{t}X^{\epsilon}_{s,u}\otimes[M^{-1}(X^{\epsilon}_{u})dW_{u}]+\int_{s}^{t}X^{\epsilon}_{s,u}\otimes[M^{-1}(X^{\epsilon}_{u})F(X^{\epsilon}_{u})du]-\epsilon\int_{s}^{t}X^{\epsilon}_{s,u}\otimes[M^{-1}(X^{\epsilon}_{u})dY^{\epsilon}_{u}]$
+$=:$ $I^{\epsilon,1}(s,t)+I^{\epsilon,2}(s,t)+I^{\epsilon,3}(s,t).$
+By Burkholder–Davis–Gundy inequality, ($\mathbf{A3}$)–($\mathbf{A4}$), Hölder inequality and Proposition 3.2 (i),
+$\mathbb{E}|I^{\epsilon,1}_{ij}(s,t)|^{p}$ $=$ $\mathbb{E}\Big|\sum_{k}\int_{s}^{t}X^{\epsilon,i}_{s,u}M^{-1}_{jk}(X^{\epsilon}_{u})dW^{k}_{u}\Big|^{p}$ (3.15)
+$\lesssim$ $\sum_{k}\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon,i}_{s,u}M^{-1}_{jk}(X^{\epsilon}_{u})|^{2}du\Big)^{p/2}$
+$\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon,i}_{s,u}|^{2}du\Big)^{p/2}$
+$\lesssim$ $(t-s)^{\frac{p}{2}-1}\int_{s}^{t}\mathbb{E}|X^{\epsilon,i}_{s,u}|^{p}du$
+$\lesssim$ $(t-s)^{\frac{p}{2}-1}\int_{s}^{t}(u-s)^{p/2}du\lesssim(t-s)^{p},$
 and
-$\Big|\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.33)
-$\lesssim$ $\Big(\int_{s}^{t}\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{p}$
-$=$ $\Big[\Big(\int_{s}^{t}\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{2}\Big]^{p/2}$
-$\lesssim$ $\Big[\int_{s}^{t}\Big(\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|du\Big)^{2}dr\Big]^{p/2},$
-so that we notice that the estimate of $J$ is identical to $I$ in view of (3.20), from which the proof is ended. ∎
+$\mathbb{E}|I^{\epsilon,2}_{ij}(s,t)|^{p}$ $=$ $\mathbb{E}\Big|\sum_{k}\int_{s}^{t}X^{\epsilon,i}_{s,u}M^{-1}_{jk}(X^{\epsilon}_{u})F_{k}(X^{\epsilon}_{u})du\Big|^{p}$ (3.16)
+$\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon,i}_{s,u}|du\Big)^{p}$
+$\leq$ $(t-s)^{p-1}\int_{s}^{t}\mathbb{E}|X^{\epsilon,i}_{s,u}|^{p}du$
+$\lesssim$ $(t-s)^{p-1}\int_{s}^{t}(u-s)^{p/2}du$
+$\lesssim$ $(t-s)^{3p/2}\lesssim(t-s)^{p}.$
+By integration-by-part formula and (1.6),
+$I^{\epsilon,3}_{ij}(s,t)$ (3.17)
+$=$ $-\epsilon\sum_{k}\int_{s}^{t}X^{\epsilon,i}_{s,u}M^{-1}_{jk}(X^{\epsilon}_{u})dY^{\epsilon,k}_{u}$
+$=$ $-\epsilon\sum_{k}\Big\{X^{\epsilon,i}_{s,t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\int_{s}^{t}Y^{\epsilon,k}_{u}d[X^{\epsilon,i}_{s,u}M^{-1}_{jk}(X^{\epsilon}_{u})]\Big\}$
+$=$ $-\epsilon\sum_{k}X^{\epsilon,i}_{s,t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}+\epsilon\sum_{k}\int_{s}^{t}Y^{\epsilon,k}_{u}M^{-1}_{jk}(X^{\epsilon}_{u})dX^{\epsilon,i}_{u}+\epsilon\sum_{k}\int_{s}^{t}Y^{\epsilon,k}_{u}X^{\epsilon,i}_{s,u}d[M^{-1}_{jk}(X^{\epsilon}_{u})]$
+$=$ $-\epsilon\sum_{k}X^{\epsilon,i}_{s,t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}+\epsilon\sum_{k}\int_{s}^{t}Y^{\epsilon,k}_{u}M^{-1}_{jk}(X^{\epsilon}_{u})dX^{\epsilon,i}_{u}+\epsilon\sum_{k,l}\int_{s}^{t}Y^{\epsilon,k}_{u}X^{\epsilon,i}_{s,u}\partial{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{u})dX^{\epsilon,l}_{u}$
+$=$ $-\epsilon\sum_{k}X^{\epsilon,i}_{s,t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}+\sum_{k}\int_{s}^{t}Y^{\epsilon,k}_{u}M^{-1}_{jk}(X^{\epsilon}_{u})Y^{\epsilon,i}_{u}du+\sum_{k,l}\int_{s}^{t}Y^{\epsilon,k}_{u}Y^{\epsilon,l}_{u}X^{\epsilon,i}_{s,u}\partial{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{u})du$
+$=:$ $J^{\epsilon,1}_{ij}(s,t)+J^{\epsilon,2}_{ij}(s,t)+J^{\epsilon,3}_{ij}(s,t).$
+By Hölder inequality, ($\mathbf{A3}$) and Proposition 3.1,
+$\mathbb{E}|J^{\epsilon,1}_{ij}(s,t)|^{p}$ $\lesssim$ $\epsilon^{p}\mathbb{E}(|X^{\epsilon,i}_{s,t}|^{p}|Y^{\epsilon,k}_{t}|^{p})$ (3.18)
+$\lesssim$ $\epsilon^{p}\sqrt{\mathbb{E}|X^{\epsilon,i}_{s,t}|^{2p}}\sqrt{\mathbb{E}|Y^{\epsilon,k}_{t}|^{2p}}$
+$\lesssim$ $\epsilon^{p}(t-s)^{p/2}\leq(t-s)^{p},$
+where we have used the dichotomy assumption in the last inequality.
+By ($\mathbf{A3}$), Hölder inequality and Proposition 3.1,
+$\mathbb{E}|J^{\epsilon,2}_{ij}(s,t)|^{p}$ $\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}|Y^{\epsilon,k}_{u}Y^{\epsilon,i}_{u}|du\Big)^{p}$ (3.19)
+$\leq$ $(t-s)^{p-1}\int_{s}^{t}\mathbb{E}|Y^{\epsilon,k}_{u}Y^{\epsilon,i}_{u}|^{p}du\lesssim(t-s)^{p}.$
+By ($\mathbf{A3}$) and Proposition 3.1,
+$\mathbb{E}|J^{\epsilon,3}_{ij}(s,t)|^{p}$ $\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}|Y^{\epsilon,k}_{u}Y^{\epsilon,l}_{u}X^{\epsilon,i}_{s,u}|du\Big)^{p}$ (3.20)
+$\leq$ $(t-s)^{p-1}\int_{s}^{t}\mathbb{E}|Y^{\epsilon,k}_{u}Y^{\epsilon,l}_{u}X^{\epsilon,i}_{s,u}|^{p}du$
+$\leq$ $(t-s)^{p-1}\int_{s}^{t}\sqrt{\mathbb{E}|Y^{\epsilon,k}_{u}Y^{\epsilon,l}_{u}|^{2p}}\sqrt{\mathbb{E}|X^{\epsilon,i}_{s,u}|^{2p}}du$
+$\lesssim$ $(t-s)^{p-1}\int_{s}^{t}(u-s)^{p/2}du$
+$\lesssim$ $(t-s)^{3p/2}\lesssim(t-s)^{p}.$
+Combining (3.17)–(3.20),
+$\mathbb{E}|I^{\epsilon,3}_{ij}(s,t)|^{p}\lesssim|t-s|^{p}.$ (3.21)
+Combining (3.14)–(3.16) and (3.21),
+$\mathbb{E}|\mathbb{X}^{\epsilon}_{s,t}|^{p}\lesssim|t-s|^{p}.$ (3.22)
 An immediate consequence of Proposition 3.2 and Kolmogorov criterion for rough paths [8, Theorem 3.1] is the following corollary.
-Corollary 3.4.
+Corollary $3.$
 Under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$), for all $1\leq p<\infty$ and $\frac{1}{3}<\alpha<\frac{1}{2},$
@@
 Remark 3.4.
-From Corollary 3.4, we see that $(X^{\epsilon})_{0<\epsilon\leq 1}$ is tight on $C([0,T];\mathbb{R}^{d}),$ which is assumed in [12]. But we do not use tightness throughout this paper.
+(i) From Corollary 3.3 and Kolmogorov tightness criterion for rough paths[8, Theorem 3.10], we see that $(\mathbf{X}^{\epsilon})_{0<\epsilon\leq 1}$ is tight in $\mathscr{C}^{0,\alpha}([0,T];\mathbb{R}^{d}).$
+[…]
+$\sup\limits_{0<\epsilon\leq 1}\mathbb{E}\sup\limits_{0\leq t\leq T}|X^{\epsilon}_{t}|^{p}<\infty,\quad p\geq 1.$ (3.23)
+Remark 3.5.
+One might expect to show Proposition 3.2 (i) by integrating factor method in classical ODE theory, that is, write
+$X^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{0}^{s}z^{\epsilon,r}_{s}dW_{r}ds+\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{0}^{s}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})drds,$
+where $z^{\epsilon,r}_{s}=\phi^{\epsilon}_{s}(\phi^{\epsilon}_{r})^{-1},\kern 5.0pt0\leq r\leq s\leq T$, and $\phi^{\epsilon}$ is the solution of the random ODE
+$\dot{\phi}^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})\phi^{\epsilon}_{t},\kern 5.0pt\phi^{\epsilon}_{0}=\mathrm{I}_{d};$
+see [27] for detailed computation.
+However, this seems incorrect, since the integrand $r\mapsto z^{\epsilon,r}_{s}$ is not $\mathcal{F}_{r}$-adapted, let alone progressively measurable, so the integrand with respect to Brownian motion is not well-defined.
+This is also pointed out by Blassel [2].
 4 Averaging Principle
@@
 $|A(x)|\leq\int_{0}^{\infty}|e^{-M^{\top}(x)t}|\cdot|B(x)|\cdot|e^{-M(x)t}|dt\leq\int_{0}^{\infty}e^{-2\lambda t}|B(x)|dt=\frac{1}{2\lambda}|B(x)|.$
-Therefore, if $B$ is of polynomial growth, then so is $A$.
+Therefore, AA has polynomial growth whenever $B$ does.
 Proof of Proposition 4.1:
@@
 so the polynomial growth property is obtained by combining (4.9), (4.13), (4.14) and Remark 4.4.
-5 Proof of the Main Theorem
-With preparation above, we prove the main theorem now.
+5 Proof of Main Theorems
+With preparation above, we prove the main theorems now.
 Proof of Theorem 2.2:
-By Corollary 3.4 and [9, Theorem A.15], it suffices to prove the pointwise convergence, that is,
+By Corollary 3.3 and interpolation theorem [9, Theorem A.15], it suffices to prove the pointwise convergence, that is,
 $X^{\epsilon}_{t}\to X_{t},\quad\text{in}\kern 5.0ptL^{p}(\Omega,\mathcal{F},P)$
@@
 for each $0\leq t\leq T.$
-Indeed, from (2.3),
-$\epsilon dY^{\epsilon}_{t}=-\frac{1}{\epsilon}M(X^{\epsilon}_{t})Y^{\epsilon}_{t}dt+F(X^{\epsilon}_{t})dt+dW_{t}=-M(X^{\epsilon}_{t})dX^{\epsilon}_{t}+F(X^{\epsilon}_{t})dt+dW_{t}.$
-Rearranging,
-$dX^{\epsilon}_{t}=M^{-1}(X^{\epsilon}_{t})dW_{t}+M^{-1}(X^{\epsilon}_{t})F(X^{\epsilon}_{t})dt-\epsilon M^{-1}(X^{\epsilon}_{t})dY^{\epsilon}_{t},$ (5.1)
-from which
+From (3.3)
 $X^{\epsilon}_{t}$ $=$ $\Big[\int_{0}^{t}M^{-1}(X^{\epsilon}_{s})dW_{s}+\int_{0}^{t}M^{-1}(X^{\epsilon}_{s})F(X^{\epsilon}_{s})ds\Big]-\epsilon\int_{0}^{t}M^{-1}(X^{\epsilon}_{s})dY^{\epsilon}_{s}$ (5.1)
@@
 $\mathbf{I}^{i}(\epsilon,t)=\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})F_{k}(X^{\epsilon}_{s})ds.$ (5.2)
-By integration-by-part formula,
-$\mathbf{J}^{i}(\epsilon,t)$ (5.4)
-$=$ $\epsilon\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})dY^{\epsilon,k}_{s}$
-$=$ $\epsilon\sum_{k}[M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\int_{0}^{t}Y^{\epsilon,k}_{s}dM^{-1}_{ik}(X^{\epsilon}_{s})]$
-$=$ $\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\epsilon\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})dX^{\epsilon,l}_{s}$
-$=$ $\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})ds.$
+From (3.8) we see that
+$\mathbf{J}^{i}(\epsilon,t)=\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})ds.$ (5.3)
 Combining (5.1), (5.2) and (5.3),
@@
 $\lesssim$ $\sum_{k}\mathbb{E}\Big(\int_{0}^{t}|M^{-1}_{ik}(X^{\epsilon}_{s})-M^{-1}_{ik}(X_{s})|^{2}ds\Big)^{p/2}$
-$\lesssim$ $\sum_{k}\mathbb{E}\Big(\int_{0}^{t}|X^{\epsilon}_{s}-X_{s}|^{2}ds\Big)^{p/2}$
-$\lesssim$ $d\mathbb{E}\int_{0}^{t}|X^{\epsilon}_{s}-X_{s}|^{p}ds$
 $\lesssim$ $\int_{0}^{t}\mathbb{E}|X^{\epsilon}_{s}-X_{s}|^{p}ds.$
-($\mathbf{A3}$) and ($\mathbf{A4}$) imply that
+Assumptions ($\mathbf{A3}$) and ($\mathbf{A4}$) imply that
 $|\mathbf{K}_{2}(\epsilon,t)|$
@@
 $=:$ $\mathbf{A}(\epsilon,t)-\mathbf{B}(\epsilon,t).$
-We explore the limit of (5.14) as $\epsilon\to 0$ componentwise.
-For $\mathbf{A}(\epsilon,t),$ by (5.11), Burkholder–Davis–Gundy inequality and dominated convergence theorem with the help of Proposition 3.1,
+We explore the limit of (5.13) as $\epsilon\to 0$ componentwise. For $\mathbf{A}(\epsilon,t),$ by (5.10), Burkholder–Davis–Gundy inequality and dominated convergence theorem with the help of (3.23),
 $\mathbf{A}^{ij}(\epsilon,t)$ (5.14)
@@
 $\to$ $\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})F_{k}(X_{s})ds,$
-where the limit is in $L^{p}(\Omega,\mathcal{F},P).$ For $\mathbf{B}(\epsilon,t),$ integrating by parts,
+where the limit is in $L^{p}(\Omega,\mathcal{F},P).$ For $\mathbf{B}(\epsilon,t),$ we see from (3.17) that
 $\mathbf{B}^{ij}(\epsilon,t)$ (5.15)
-$=$ $\epsilon\Big(\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})dY^{\epsilon}_{s}]\Big)_{ij}$
-$=$ $\epsilon\int_{0}^{t}X^{\epsilon,i}_{s}[M^{-1}(X^{\epsilon}_{s})dY^{\epsilon}_{s}]^{j}$
-$=$ $\epsilon\sum_{k}\int_{0}^{t}X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})dY^{\epsilon,k}_{s}$
-$=$ $\epsilon\sum_{k}\Big(X^{\epsilon,i}_{t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\int_{0}^{t}Y^{\epsilon,k}_{s}d[X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})]\Big)$
-$=:$ $\mathbf{B}^{ij}_{1}(\epsilon,t)-\mathbf{B}^{ij}_{2}(\epsilon,t).$
-By Proposition 3.1 and ($\mathbf{A3}$), $(X^{\epsilon,i}_{t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t})_{0<\epsilon\leq 1}$ is bounded in $L^{p}(\Omega,\mathcal{F},P)$ for each $t\in[0,T],$ so that
+$=$ $\epsilon\sum_{k}X^{\epsilon,i}_{t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})ds-\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}X^{\epsilon,i}_{s}\partial_{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{s})ds$
+$=:$ $\mathbf{B}^{ij}_{1}(\epsilon,t)-\mathbf{B}^{ij}_{2}(\epsilon,t)-\mathbf{B}^{ij}_{3}(\epsilon,t).$
+By Proposition 3.1, (3.23) and ($\mathbf{A3}$), $(X^{\epsilon,i}_{t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t})_{0<\epsilon\leq 1}$ is bounded in $L^{p}(\Omega,\mathcal{F},P)$ for each $t\in[0,T],$ so that
 $\mathbf{B}_{1}^{ij}(\epsilon,t)\to 0\quad\text{in}\quad L^{p}(\Omega,\mathcal{F},P).$ (5.16)
-On the other hand, notice that
-$\mathbf{B}_{2}^{ij}(\epsilon,t)$ (5.18)
-$=$ $\epsilon\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}d[X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})]$
-$=$ $\epsilon\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})dX^{\epsilon,i}_{s}+\epsilon\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}X^{\epsilon,i}_{s}\sum_{l}[\partial_{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{s})dX^{\epsilon,l}_{s}]$
-$=$ $\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})ds+\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}X^{\epsilon,i}_{s}\partial_{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{s})ds$
-$=:$ $\mathbf{B}_{21}^{ij}(\epsilon,t)+\mathbf{B}_{22}^{ij}(\epsilon,t).$
 By (5.12),
-$\mathbf{B}_{21}^{ij}(\epsilon,t)\to\sum_{k}\int_{0}^{t}M^{-1}_{jk}(X_{s})J_{ik}(X_{s})ds,$ (5.19)
+$\mathbf{B}_{2}^{ij}(\epsilon,t)\to\sum_{k}\int_{0}^{t}M^{-1}_{jk}(X_{s})J_{ik}(X_{s})ds,$ (5.17)
 and
-$\mathbf{B}_{22}^{ij}(\epsilon,t)\to\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds,$ (5.20)
+$\mathbf{B}_{3}^{ij}(\epsilon,t)\to\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds,$ (5.18)
 where the two limits above are both in $L^{p}(\Omega,\mathcal{F},P).$ Combining (5.14), (5.16), (5.17) and (5.18), we conclude that
@@
 $\int_{0}^{t}X^{i}_{s}dX^{j}_{s}$
-$=$ $\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{ik}(X_{s})F_{k}(X_{s})ds+\sum_{k}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s},$
+$=$ $\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})F_{k}(X_{s})ds+\sum_{k}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s},$
 and
 $[X^{i},X^{j}]_{t}=\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds.$
-However, multipling both sides of (1.3) by $M^{-1}$ on the left and $(M^{-1})^{\top}$ on the right and then writing it componentwise, we see that
+However, multipling both sides of (1.9) by $M^{-1}$ on the left and $(M^{-1})^{\top}$ on the right handside and then writing it componentwise, we see that
 $\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds=\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})J_{kj}(X_{s})+J_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds.$
@@
 $\int_{0}^{t}X^{i}_{s}\circ dX^{j}_{s}$
-$=$ $\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{ik}(X_{s})F_{k}(X_{s})ds$
+$=$ $\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})F_{k}(X_{s})ds$
 $+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s}+\frac{1}{2}\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})J_{kj}(X_{s})+J_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds.$
@@
 $\Big(\int_{0}^{t}X_{s}\otimes\circ dX_{s}\Big)_{ij}+\frac{1}{2}\Big(\int_{0}^{t}J(X_{s})(M^{-1})^{\top}(X_{s})-M^{-1}(X_{s})J(X_{s})ds\Big)_{ij}$ (5.20)
-$=$ $\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{ik}(X_{s})F_{k}(X_{s})ds$
+$=$ $\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})F_{k}(X_{s})ds$
 $+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}J_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds,$
-which coincides with (5.21).
+which coincides with (5.19).∎
+Before demonstrating Theorem 2.4, we discuss a slightly more general result about transferring RDE to SDE. Consider the following RDE
+$dZ_{t}=b(Z_{t})dt+\sigma(Z_{t})d\mathbf{Y}_{t},$
+where $\mathbf{Y}=(Y,\mathbb{Y})$ is a $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$-valued random element, $Y$ is a continuous semimartingale,
+[…]
+the later integral being Itô integral, and $a$ is a $C([0,T];\mathbb{R}^{d\times d})$-valued random element. Then for almost every sample path, we have
+$Z_{t}-Z_{s}$ (5.21)
+$=$ $b(Z_{s})(t-s)+\sigma(Z_{s})Y_{s,t}+D\sigma(Z_{s})\sigma(Z_{s})\mathbb{Y}_{s,t}+R_{s,t},$
+with $\lim\limits_{|\mathcal{P}|\to 0}\sup\limits_{\mathcal{P}\subset[0,T]}\sum\limits_{[s,t]\in\mathcal{P}}|R_{s,t}|=0,$ $P$-a.s. . But
+$(D\sigma(Z_{s})\sigma(Z_{s})\mathbb{Y}_{s,t})_{i}$ (5.22)
+$=$ $\sum_{j,k,l}\partial_{z_{k}}\sigma_{ij}(Z_{s})\sigma_{kl}(Z_{s})\mathbb{Y}^{lj}_{s,t}$
+[…]
+[…]
+where $\lim\limits_{|\mathcal{P}|\to 0}\sup\limits_{\mathcal{P}\subset[0,T]}\sum\limits_{[s,t]\in\mathcal{P}}|\hat{R}_{s,t}|=0,$ $P$-a.s. . Combining (5.21) and (5.22),
+$(Z_{t}-Z_{s})_{i}$
+$=$ $[b_{i}(Z_{s})+\sum_{j,k,l}\partial_{z_{k}}\sigma_{ij}(Z_{s})\sigma_{kl}(Z_{s})a^{lj}_{s}](t-s)$
+[…]
+where $\lim\limits_{|\mathcal{P}|\to 0}\sup\limits_{\mathcal{P}\subset[0,T]}\sum\limits_{[s,t]\in\mathcal{P}}|\tilde{R}^{i}_{s,t}|=0,$ $P$-a.s.
+.
+By equivalence theorem of RDE and SDE[10, Proposition 6.9], for each $1\leq i\leq d,$
+$dZ^{i}_{t}=[b_{i}(Z_{t})+\sum\limits_{j,k,l}\partial_{z_{k}}\sigma_{ij}(Z_{t})\sigma_{kl}(Z_{t})a^{lj}_{t}]dt+\sum_{j}\sigma_{ij}(Z_{t})dY^{j}_{t}.$ (5.23)
+Proof of Theorem 2.4:
+By Theorem 2.2 and a probabilistic version of continuity of Itô–Lyons map[10, Theorem 6.1], $Z^{\epsilon}\to Z$ in $C^{\alpha}([0,T];\mathbb{R}^{d})$ in probability, with
+$dZ_{t}=b(Z_{t})dt+\sigma(Z_{t})d\mathbf{X}_{t}\,.$
+Now we transfer this RDE to SDE. By the same argument leading to (5.20), we have
+$\mathbb{X}^{ij}_{s,t}$ (5.24)
+$=$ $\Big(\int_{s}^{t}X_{u}\otimes\circ dX_{u}\Big)_{ij}+\frac{1}{2}\Big(\int_{s}^{t}J(X_{u})(M^{-1})^{\top}(X_{u})-M^{-1}(X_{u})J(X_{u})du\Big)_{ij}$
+$=$ $\sum_{k,l}\int_{s}^{t}X^{i}_{u}S_{j}(X_{u})du+\sum_{k}\int_{s}^{t}X^{i}_{u}M^{-1}_{jk}(X_{u})F_{k}(X_{u})du$
+$+\sum_{k}\int_{s}^{t}X^{i}_{u}M^{-1}_{jk}(X_{u})dW^{k}_{u}+\sum_{k}\int_{s}^{t}J_{ik}(X_{u})M^{-1}_{jk}(X_{u})du$
+[…]
+Applying (5.23),
+$dZ^{i}_{t}$
+$=$ $[b_{i}(Z_{t})+\sum\limits_{j,k,l,k^{\prime}}\partial_{z_{k}}\sigma_{ij}(Z_{t})\sigma_{kl}(Z_{t})J_{lk^{\prime}}(X_{t})M^{-1}_{jk^{\prime}}(X_{t})]dt+\sum_{j}\sigma_{ij}(Z_{t})dY^{j}_{t}$
+$=$ $[b_{i}(Z_{t})+\sum\limits_{j,k,l,k^{\prime}}\partial_{z_{k}}\sigma_{ij}(Z_{t})\sigma_{kl}(Z_{t})J_{lk^{\prime}}(X_{t})M^{-1}_{jk^{\prime}}(X_{t})]dt$
+$+[\sum_{j}\sigma_{ij}(Z_{t})S_{j}(X_{t})+\sum_{j,k}\sigma_{ij}(Z_{t})M^{-1}_{jk}(X_{t})F_{k}(X_{t})]dt+\sum_{j,k}\sigma_{ij}(Z_{t})M^{-1}_{jk}(X_{t})dW^{k}_{t}.$